### v1

# Unified interpretation of 95 GeV di-photon and di-tau Excesses in the Georgi-Machacek Model

Abstract
We revisit the 95 GeV excesses in the $\gamma\gamma$ and $\tau\tau$ channels in the Georgi–Machacek model, where a single light $CP$-even custodial singlet $H$ can account for both hints. Using one-loop renormalization group-improved effective potential and positive-definiteness conditions for vacuum stability, together with perturbative unitarity, electroweak precision tests, $B$-physics, and Higgs data, we identify a narrow but viable parameter region. In this region, charged and doubly charged scalars enhance $H\to\gamma\gamma$, while the fermionic rescaling controls $H\to\tau\tau$, yielding combined signal strengths compatible at the $2\sigma$ level. The scenario predicts characteristic patterns in the singlet mixing and triplet VEV and is highly testable at the HL-LHC and future lepton colliders via precision $\kappa_{V}$ measurements and direct exotic searches.
1 Introduction
The discovery of the 125 GeV Higgs boson by the ATLAS ATLAS:2012ae and CMS CMS:2012zhx collaborations in 2012 confirms the Standard Model (SM) framework of electroweak symmetry breaking (EWSB). Nevertheless, theoretical motivations and the sizable uncertainties in current Higgs data fits still leave ample room for an extended Higgs sector ATLAS:2019nkf . Consequently, the search for additional scalar fields remains a crucial objective.
[…]
The Georgi-Machacek (GM) model Georgi:1985nv ; Chanowitz:1985ug ; Gunion:1989ci ; Gunion:1990dt ; Gunion:1989we ; Haber:1999zh ; Godfrey:2010qb ; Logan:2010en ; Low:2012rj ; Killick:2013mya ; Efrati:2014uta ; Degrande:2015xnm extends the SM by introducing a complex ${\rm SU(2)}_{L}$ triplet scalar $\chi$ with hypercharge +1+1 and a real ${\rm SU(2)}_{L}$ triplet scalar $\xi$ with hypercharge 0. The field content is arranged such that custodial symmetry is preserved at tree level after EWSB, provided the triplet vacuum expectation values (VEVs) are aligned, $v_{\chi}=v_{\xi}\equiv v_{\Delta}$. This alignment permits $v_{\Delta}$ values up to several tens GeV, significantly larger than in generic triplet extensions, while maintaining $\rho_{\rm tree}=1$. The model predicts nine additional scalar states, including a characteristic doubly charged Higgs boson, resulting in a rich phenomenology and distinctive collider signatures Chiang:2012cn ; Kanemura:2013mc ; Englert:2013zpa ; Englert:2013wga ; Godunov:2014waa ; Chiang:2015kka ; Chiang:2015amq ; Yu:2016skv ; Cao:2016rht ; Chang:2017niy ; Zhang:2017och ; Li:2017daq ; Sun:2017mue ; Azuelos:2017dqw ; Cao:2018xwy ; Yang:2018pzt ; Ghosh:2019qie ; Ismail:2020zoz ; Zhu:2020bew ; Wang:2022okq ; Ghosh:2022wbe ; deLima:2022yvn ; Ahriche:2022aoj ; Ghosh:2022bvz ; Chakraborti:2023mya ; Ghosh:2023izq ; Bhattacharya:2024gyh ; Lu:2024ade ; Ghosh:2025htt ; Ashanujjaman:2025puu ; Ghosh:2025hyy ; Chiang:2025lab ; Chiang:2018cgb . Notably, the GM model allows the coupling of the SM-like Higgs boson to weak gauge bosons to exceed the SM value Falkowski:2012vh ; Chang:2012gn ; Chiang:2013rua .
The extended scalar sector improves vacuum stability and can induce strong first-order electroweak phase transitions (EWPT), potentially generating gravitational wave signals accessible to future experiments Chiang:2014hia ; Chen:2022zsh ; Zhou:2018zli ; Bian:2019bsn ; Lu:2025vif . If lepton number violation is present, triplet-lepton interactions can generate neutrino masses Chen:2020ark ; Giarnetti:2023dcr ; Giarnetti:2023osf ; Du:2022brr . The GM framework has also been employed to address the $W$-mass anomaly reported by CDF-II Du:2022brr ; Mondal:2022xdy ; Chen:2022ocr ; Ghosh:2022zqs , especially when custodial symmetry breaking effects are included Blasi:2017xmc ; Keeshan:2018ypw ; Song:2022jns . Furthermore, the light custodial singlet scalar of the GM model constitutes a natural candidate to interpret the excesses observed near 95 GeV Ahriche:2023wkj ; Chen:2023bqr ; Du:2025eop ; Ghosh:2025ebc ; Mondal:2025tzi . A unified explanation for the excesses observed in the $\gamma\gamma$, $\tau\tau$, and $b\bar{b}$ channels requires a near-degenerate CPCP-even/CPCP-odd ‘twin-peak’ structure Chen:2023bqr ; with only a single light $CP$-even state, the $\tau\tau$ excess is in tension with the $\gamma\gamma$ and $b\bar{b}$ channels Ahriche:2023wkj ; Chen:2023bqr ; Du:2025eop . Recent analyses, however, strongly disfavor the LEP $b\bar{b}$ excess as evidence for a new scalar Janot:2024ryq , motivating a focus on the LHC-driven $\gamma\gamma$ and $\tau\tau$ channels. In this context, the newly proposed positive-definiteness vacuum stability conditions based on one-loop renormalization group-improved (RG-improved) effective potential Du:2024mry expand the viable parameter space beyond that allowed by traditional bounded-from-below requirements. This enables a unified interpretation of both excesses within a single $CP$-even Higgs boson scenario.
In this work, we focus on the unified interpretation of $\gamma\gamma$ and $\tau\tau$ resonance signals within the GM model. This work is organised as follows. In Sec. 2, we provide a brief review of the GM model. A short discussion about the di-photon and di-tau excess signals at a mass of 95 GeV is presented in Sec. 3. Theoretical and experimental constraints relevant to realistic research within the GM model are detailed in Sec. 4. Numerical results are displayed in Sec. 5. Finally, we present our conclusions in Sec. 6.
2 The Georgi-Machacek model
The GM model extends the SM scalar sector by introducing additional triplet fields. It contains the familiar ${\rm SU(2)}_{L}$ doublet $({\phi}^{+},~{\phi}^{0})$ with hypercharge $1/2$, a complex triplet $({\chi}^{++},~{\chi}^{+},~{\chi}^{0})$ with hypercharge +1+1, and a real triplet $({\xi}^{+},~{\xi}^{0},~-{\xi}^{+*})$ with zero hypercharge. This field structure enables custodial ${\rm SU(2)}_{V}$ symmetry at tree level and leads to distinctive scalar phenomenology.
To clarify custodial symmetry, the doublet and triplets are arranged in matrices $\Phi$ and $\Delta$, respectively:
$\Phi=\begin{pmatrix}\phi^{0*}&\phi^{+}\\ \phi^{-}&\phi^{0}\end{pmatrix},\quad\Delta=\begin{pmatrix}\chi^{0*}&\xi^{+}&\chi^{++}\\ \chi^{-}&\xi^{0}&\chi^{+}\\ \chi^{--}&\xi^{-}&\chi^{0}\end{pmatrix}$. (1)
Then the most general scalar potential for the GM scalar sector, constructed to be invariant under a global ${\rm SU(2)}_{L}\times{\rm SU(2)}_{R}$ symmetry, hence also under its gauged subgroup ${\rm SU(2)}_{L}\times{\rm U(1)}_{Y}$, reads
$\begin{split}V(\Phi,\Delta)=&\frac{1}{2}m_{\Phi}^{2}{\rm tr}\left[\Phi^{\dagger}\Phi\right]+\frac{1}{2}m_{\Delta}^{2}{\rm tr}\left[\Delta^{\dagger}\Delta\right]\\ &+\lambda_{1}\left({\rm tr}\left[\Phi^{\dagger}\Phi\right]\right)^{2}+\lambda_{2}\left({\rm tr}\left[\Delta^{\dagger}\Delta\right]\right)^{2}+\lambda_{3}{\rm tr}\left[\left(\Delta^{\dagger}\Delta\right)^{2}\right]\\ &+\lambda_{4}{\rm tr}\left[\Phi^{\dagger}\Phi\right]{\rm tr}\left[\Delta^{\dagger}\Delta\right]+\lambda_{5}{\rm tr}\left[\Phi^{\dagger}\frac{\sigma^{a}}{2}\Phi\frac{\sigma^{b}}{2}\right]{\rm tr}\left[\Delta^{\dagger}T^{a}\Delta T^{b}\right]\\ &+\mu_{1}{\rm tr}\left[\Phi^{\dagger}\frac{\sigma^{a}}{2}\Phi\frac{\sigma^{b}}{2}\right](P^{\dagger}\Delta P)_{ab}+\mu_{2}{\rm tr}\left[\Delta^{\dagger}T^{a}\Delta T^{b}\right](P^{\dagger}\Delta P)_{ab},\end{split}$ (2)
Here, $\sigma^{i}$ are Pauli matrices, and $T^{i}$ are the $\rm SU(2)$} generators in the triplet representation:
$T^{1}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&1&0\\ 1&0&1\\ 0&1&0\end{pmatrix},\quad T^{2}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&-\mathrm{i}&0\\ \mathrm{i}&0&-\mathrm{i}\\ 0&\mathrm{i}&0\end{pmatrix},\quad T^{3}=\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&-1\end{pmatrix}$. (3)
The $P$ matrix rotates the triplet fields $\Delta$ to the Cartesian basis
$P=$ $\frac{1}{\sqrt{2}}\begin{pmatrix}-1&\mathrm{i}&0\\ 0&0&\sqrt{2}\\ 1&\mathrm{i}&0\end{pmatrix}$. (4)
It should be noted that the last two terms in Eq. (2) vanish under a $\mathbb{Z}_{2}$ symmetry. However, we retain them in this work.
Under ${\rm SU(2)}_{L}\times{\rm SU(2)}_{R}$, the fields $\Phi$ and $\Delta$ transform as
$\Phi\to U_{L}\Phi U_{R}^{\dagger},\quad\Delta\to U_{L}\Delta U_{R}^{\dagger}.$ (5)
Here, $U_{L,R}$ denotes $\exp(\mathrm{i}\theta^{a}_{L,R}\sigma^{a})$ for the doublet and $\exp(\mathrm{i}\theta^{a}_{L,R}T^{a})$ for the triplet, respectively. The SM electroweak gauge group ${\rm SU(2)}_{L}\times{\rm U(1)}_{Y}$ is embedded in the global ${\rm SU(2)}_{L}\times{\rm SU(2)}_{R}$ symmetry, with ${\rm SU(2)}_{L}$ identified by the generators $T_{L}^{i}$ and the hypercharge ${\rm U(1)}_{Y}$ associated with the generator $T_{R}^{3}$.
After the EWSB, the neutral scalar fields can be parameterized as
$\begin{split}{\xi}^{0}&\to v_{\xi}+h_{\xi},\\ {\phi}^{0}&\to\frac{1}{\sqrt{2}}\left(v_{\phi}+h_{\phi}+ia_{\phi}\right)\\ {\chi}^{0}&\to v_{\chi}+\frac{1}{\sqrt{2}}\left(h_{\chi}+ia_{\chi}\right).\end{split}$ (6)
Here, the parameters $v_{\xi}$, $v_{\phi}$ and $v_{\chi}$ represent the VEVs of the corresponding neutral components of these scalar fields. Then the tree level $\rho$ parameter of the GM model can be expressed as
$\rho_{\rm tree}\equiv\frac{m^{2}_{W}}{m_{Z}^{2}{\cos^{2}{\theta_{W}}}}=\frac{v_{\phi}^{2}+4v_{\chi}^{2}+4v_{\xi}^{2}}{v_{\phi}^{2}+8v_{\chi}^{2}}.$ (7)
Therefore, when the triplet VEVs are aligned, $v_{\chi}=v_{\xi}\equiv v_{\Delta}$, the vacuum remains invariant only under the diagonal (custodial) subgroup ${\rm SU(2)}_{V}\subset{\rm SU(2)}_{L}\times{\rm SU(2)}_{R}$ (i.e., when $U_{L}=U_{R}$). As a result, the tree-level $\rho$ parameter automatically satisfies $\rho_{\rm tree}=1$. In this alignment limit, the VEVs of scalar field take the form
$\langle\Phi\rangle=\frac{v_{\phi}}{\sqrt{2}}{\bf I}_{2\times 2},\quad\text{and}\quad\langle\Delta\rangle=v_{\Delta}{\bf I}_{3\times 3}.$ (8)
And the condition for EWSB is given as
$v^{2}=v_{\phi}^{2}+8v_{\Delta}^{2}\simeq\left(246~{\rm GeV}\right)^{2}.$ (9)
The physical scalars are then decomposed into irreducible representations according to the custodial ${\rm SU(2)}_{V}$ symmetry. The doublet $\Phi$ (interpreted as a ${\bf 2}\otimes{\bf 2}$ representation) decomposes into ${\bf 1}\oplus{\bf 3}$, while the triplet $\Delta$ (as a ${\bf 3}\otimes{\bf 3}$ representation) decomposes into ${\bf 1}\oplus{\bf 3}\oplus{\bf 5}$. One triplet, $G_{3}$, provides the Nambu-Goldstone bosons eaten by $W$ and $Z$ bosons
$G_{3}^{\pm}=\cos{\theta}\phi^{\pm}+\frac{\sin{\theta}}{\sqrt{2}}\left(\chi^{\pm}+\xi^{\pm}\right),\quad G_{3}^{0}=\cos{\theta}{a_{\phi}}+\sin{\theta}a_{\chi},$ (10)
where the mixing parameter $\theta$ is defined by
$\tan{\theta}\equiv 2\sqrt{2}v_{\Delta}/v_{\phi}.$ (11)
The singlet $H_{1}^{0}$ (originating from $\Delta$ field), the triplet ($H_{3}^{\pm}$ and $H_{3}^{0}$), and the quintuplet ($H_{5}^{\pm\pm}$, $H_{5}^{\pm}$ and $H_{5}^{0}$) are given by
$\begin{split}H^{0}_{1}&=\sqrt{\frac{2}{3}}h_{\chi}+\sqrt{\frac{1}{3}}h_{\xi},\\ H_{3}^{\pm}&=-\sin{\theta}\phi^{\pm}+\frac{\cos{\theta}}{\sqrt{2}}\left(\chi^{\pm}+\xi^{\pm}\right),\quad H_{3}^{0}=\cos{\theta}a_{\chi}-\sin{\theta}a_{\phi},\\ H_{5}^{\pm\pm}&=\chi^{\pm\pm},\quad H_{5}^{\pm}=\frac{1}{\sqrt{2}}\left(\chi^{\pm}-\xi^{\pm}\right),\quad H_{5}^{0}=\sqrt{\frac{2}{3}}h_{\xi}-\sqrt{\frac{1}{3}}h_{\chi}.\end{split}$ (12)
In the basis of $(h_{\phi},H_{1}^{0},H_{5}^{0})$, the mass matrix of $CP$-even Higgs states is given by
$M^{2}_{CP\text{-even}}=\begin{pmatrix}M_{11}^{2}&M_{12}^{2}&0\\ M_{12}^{2}&M_{22}^{2}&0\\ 0&0&m_{H_{5}}^{2}\end{pmatrix}$, (13)
where the elements of the $2\times 2$ part are
$\begin{split}{M}_{11}^{2}&=8\lambda_{1}v_{\phi}^{2},\\ {M}_{12}^{2}={M}_{21}^{2}&=\frac{\sqrt{3}}{2}v_{\phi}\left(4\left(2\lambda_{2}+\lambda_{5}\right)v_{\Delta}+\mu_{1}\right),\\ {M}_{22}^{2}&=8\left(\lambda_{3}+4\lambda_{4}\right)v_{\Delta}^{2}-\frac{v_{\phi}^{2}}{4v_{\Delta}}\mu_{1}+6v_{\Delta}\mu_{2}.\end{split}$ (14)
The mass eigenstates $h$ and $H$ are defined by
$h=\cos{\alpha}h_{\phi}-\sin{\alpha}{H^{0}_{1}},\quad H=\sin{\alpha}h_{\phi}+\cos{\alpha}{H^{0}_{1}},$ (15)
where the mixing angle $\alpha$ is determined by
$\tan{2\alpha}=\frac{2M_{12}^{2}}{M_{22}^{2}-M_{11}^{2}},\quad\alpha\in\left(0,\pi\right)$ (16)
The masses of all Higgs mass eigenstates can then be expressed as
$\begin{split}m_{h}^{2}&=M_{11}^{2}\cos^{2}{\alpha}+M_{22}^{2}\sin^{2}{\alpha}+M_{12}^{2}\sin{2\alpha},\\ m_{H}^{2}&=M_{11}^{2}\sin^{2}{\alpha}+M_{22}^{2}\cos^{2}{\alpha}-M_{12}^{2}\sin{2\alpha},\\ m_{H_{3}}^{2}\equiv m_{H_{3}^{\pm}}^{2}=m_{H_{3}^{0}}^{2}&=-\frac{v^{2}}{4v_{\Delta}}\mu_{1}-\frac{1}{2}\lambda_{5}v^{2},\\ m_{H_{5}}^{2}\equiv m_{H_{5}^{\pm\pm}}^{2}=m_{H_{5}^{\pm}}^{2}=m_{H_{5}^{0}}^{2}&=-\frac{v_{\phi}^{2}}{4v_{\Delta}}\mu_{1}-12v_{\Delta}\mu_{2}+8\lambda_{3}v_{\Delta}^{2}-\frac{3}{2}\lambda_{5}v_{\phi}^{2}.\end{split}$ (17)
By applying the tadpole conditions,
$\frac{\partial V}{\partial h_{\phi}}\bigg|_{0}=0,\quad\frac{\partial V}{\partial h_{\xi}}\bigg|_{0}=0,\quad\frac{\partial V}{\partial h_{\chi}}\bigg|_{0}=0,$ (18)
the mass parameters $m_{\phi}^{2}$ and $m_{\Delta}^{2}$ in Eq. (2) can be eliminated as follows:
$\begin{split}m_{\phi}^{2}&=-4\lambda_{1}v_{\phi}^{2}-\left(\frac{3}{2}\mu_{1}+6\lambda_{4}+3\lambda_{5}\right)v_{\Delta}^{2},\\ m_{\Delta}^{2}&=-(12\lambda_{2}+4\lambda_{3})v_{\Delta}^{2}-\left(\frac{1}{4v_{\Delta}}\mu_{1}+2\lambda_{4}+\lambda_{5}\right)v_{\phi}^{2}-6v_{\Delta}\mu_{2}.\end{split}$ (19)
In summary, the free parameters of the GM model consist of five dimensionless couplings $\lambda_{1,2,3,4,5}$, two dimensional parameters $\mu_{1,2}$, and one mixing parameter $\theta$. In practice, we adopt the following set of free parameters in this work:
$\lambda_{2,3,4,5},\quad m_{h}^{\rm In},\quad m_{H}^{\rm In},\quad\alpha,\quad\sin{\theta},$ (20)
where $\mu_{1,2}$ are substituted by $m_{h,H}^{\rm In}$, and $\lambda_{1}$ is traded for $\alpha$. In the following sections, $m_{h,H}$ without a superscript refers to the physical pole mass as calculated by the spectrum calculator.
3 95 GeV excesses in the GM Model
In the setup introduced in Sec. 2, the GM model contains two light singlet scalar fields. In this work, the heavier state, denoted as $h$, is identified with the $125~\mathrm{GeV}$ SM-like Higgs boson. The other state, labeled as $H$, serves as a candidate to explain the observed excesses around $95~\mathrm{GeV}$. The couplings of these Higgs states to SM fermions ($f$) and weak gauge bosons ($V=W,Z$) are modified as follows:
$\begin{split}g_{hff}&=\kappa_{ff}^{h}\times g_{hff}^{\rm SM},\quad g_{hVV}=\kappa_{VV}^{h}\times g_{hVV}^{\rm SM},\\ g_{Hff}&=\kappa_{ff}^{H}\times g_{hff}^{\rm SM},\quad g_{HVV}=\kappa_{VV}^{h}\times g_{hVV}^{\rm SM},\end{split}$ (21)
with the effective Higgs couplings $\kappa$ in the lowest order
[…]
Using this framework, the GM prediction of the di-photon signal strength normalised to its SM prediction $\mu_{\gamma\gamma}$ are given by
$\begin{split}\mu_{\gamma\gamma}&\equiv\frac{\sigma_{\rm GM}\left(pp\to H_{95}\to\gamma\gamma\right)}{\sigma_{\rm SM}\left(pp\to h_{95}\to\gamma\gamma\right)}\\ &\simeq\frac{\sigma_{\rm GM}\left(gg\to H_{95}\right)}{\sigma_{\rm SM}\left(gg\to h_{95}\right)}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to\gamma\gamma\right)}{\mathcal{B}_{\rm SM}\left(h_{95}\to\gamma\gamma\right)}\\ &=\left|\kappa_{ff}^{H}\right|^{2}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to\gamma\gamma\right)}{1.39\times 10^{-3}},\end{split}$ (23)
where the mass of the Higgs boson is fixed around 95 GeV (denoted by $H_{95}$ and $h_{95}$), and the subscript $\rm GM$ ($\rm SM$) denotes the predictions of the GM model (SM) on the inclusive production rate of $H_{95}$($h_{95}$). The specific value $1.39\times 10^{-3}$ is the SM Higgs branching ratio at 95 GeV calculated by LHC Higgs Cross Section Working Group LHCHiggsCrossSectionWorkingGroup:2013rie . The second line approximation arises because the production of the SM Higgs boson is dominated by the gluon-gluon fusion process, which accounts for approximately 89% of the total Higgs production at a mass of 95 GeV LHCHiggsCrossSectionWorkingGroup:2013rie . The third line rises because the gluon-gluon fusion process ($gg\to h$) is mediated by a colored-fermion loop. As a result, the effective coupling $\kappa^{H}_{gg}$ equals $\kappa^{H}_{ff}$, since the fermionic couplings of $H$ scale universally as $\kappa_{ff}^{H}$ in lowest order, as shown in Eq. (22). Similarly, the GM prediction of the $\tau\tau$ and $bb$ channel signal strengths are given by
$\begin{split}\mu_{bb}&\equiv\frac{\sigma_{\rm GM}\left(e^{+}e^{-}\to ZH_{95}(H_{95}\to bb)\right)}{\sigma_{\rm SM}\left(e^{+}e^{-}\to Zh_{95}(h_{95}\to\gamma\gamma\right)}\\ &=\left|\kappa_{VV}^{H}\right|^{2}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to bb\right)}{0.801},\\ \mu_{\tau\tau}&\equiv\frac{\sigma_{\rm GM}\left(pp\to H_{95}\to\tau\tau\right)}{\sigma_{\rm SM}\left(pp\to h_{95}\to\tau\tau\right)}\\ &=\left|\kappa_{ff}^{H}\right|^{2}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to\tau\tau\right)}{8.32\times 10^{-2}}.\end{split}$ (24)
From above definitions, all strengths are expressed in term of the ratio $\sigma_{\rm GM}/\sigma_{\rm SM}$, so the QCD correction to the numerator and the denominator will cancel. This is beneficial to reduce the theoretical uncertainty in predicting the strengths.
Now, we scrutinize the branching ratios. The partial decay width of the lighter scalar $H$ into SM final states can be expressed as
$\Gamma(H\to XX)=|\kappa^{H}_{XX}|^{2}\times\Gamma_{\rm SM}(h\to XX)\big|_{m_{h}=m_{H}},$ (25)
where $\Gamma_{\rm SM}(h\to XX)\big|_{m_{h}=m_{H}}$ is the SM Higgs partial decay width estimated at mass of $m_{H}$. Therefore, the total decay width of custodial singlet state $H$ can be written as
$\Gamma^{\rm tot}_{H}=\Gamma^{\rm SM}_{h}\sum_{X}\left|\kappa^{H}_{XX}\right|^{2}\mathcal{B}_{\rm SM}\left(h\to XX\right),$ (26)
where the SM Higgs total decay width $\Gamma^{\rm SM}_{h}$ and the branching rations $\mathcal{B}_{\rm SM}\left(h\to XX\right)$ estimated at $m_{H}$. The coupling between neutral scalar $H$ or $h$ and photons are induced by charged particle loops. In SM, the Higgs boson decay into two photons is mediated by $W$ boson and heavy charged fermion (top quark dominated) loops. While in GM model, the heavy charged Higgs states $S_{i}$ (including $H_{5}^{\pm\pm}$, $H_{5}^{\pm}$ and $H_{3}^{\pm}$) loops can significantly enhance the partial decay width $\Gamma(H\to\gamma\gamma)$. The corresponding effective coupling $\kappa^{H}_{\gamma\gamma}$ can be expressed as
$\small\kappa^{H}_{\gamma\gamma}=\frac{\sum_{i}\dfrac{g_{HS_{i}S_{i}}}{2m_{S_{i}}^{2}}vQ_{S_{i}}^{2}A_{0}\left(\frac{m_{H}^{2}}{4m_{S_{i}}^{2}}\right)+\kappa_{VV}^{H}A_{1}\left(\frac{m_{H}^{2}}{4m_{W}^{2}}\right)+\kappa_{ff}^{H}~\dfrac{4}{3}A_{1/2}\left(\frac{m_{H}^{2}}{4m_{t}^{2}}\right)}{A_{1}\left(\frac{m_{H}^{2}}{4m_{W}^{2}}\right)+\dfrac{4}{3}A_{1/2}\left(\frac{m_{H}^{2}}{4m_{t}^{2}}\right)},$ (27)
where $g_{HS_{i}S_{i}}$ are the scalar trilinear couplings of $H$ to the charged states $S_{i}$, and complex factors $A_{0,1/2,1}$ can be found in the literature Djouadi:2005gi ; Djouadi:2005gj ; Gunion:1989we . For an $H$ boson with mass $95~{\rm GeV}$, $A_{1}\simeq-7.6$ and $A_{1/2}\approx 4/3$. The factor $A_{0}$ reaches its maximum value of $1.5+i$ at $m_{H}=2m_{S_{i}}$, while in the limit of very heavy loop charged scalars, $A_{0}(0)\simeq 1/3$.
Then the signal strength in Eq. (23) and Eq. (24) can be rewriten as:
$\begin{split}\mu_{\gamma\gamma}&=\left|\kappa^{H}_{ff}\right|^{2}\left|\kappa^{H}_{\gamma\gamma}\right|^{2}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}}\\ \mu_{bb}&=\left|\kappa^{H}_{VV}\right|^{2}\left|\kappa^{H}_{ff}\right|^{2}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}},\\ \mu_{\tau\tau}&=\left|\kappa^{H}_{ff}\right|^{4}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}}.\end{split}$ (28)
It is clear that all three signal strengths share a common factor, $\left|\kappa^{H}_{ff}\right|^{2}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}}$. This introduces some tensions in attributing the observed excesses to a single source. While charged scalar states in GM models can provide a positive contributions to interpreting the $\gamma\gamma$ excess, especially the doubly charged $H_{5}^{\pm\pm}$ benefiting from $Q_{S}^{2}=4$. Ref. Chen:2023bqr investigate the different possibility of $H$ explaining $\gamma\gamma$, $\gamma\gamma+bb$ and $\gamma\gamma+bb+\tau\tau$. However, recent dedicated reanalyses argue that this feature is unlikely to originate from a genuine new scalar resonance Janot:2024ryq . Given the inferior dijet mass resolution at LEP and the updated conclusions, we do not attempt to accommodate the $b\bar{b}$ hint in this work. Instead, we focus on the two LHC-driven channels ($\gamma\gamma$ and $\tau\tau$) where the experimental systematics and resolutions are better controlled in the relevant mass window. Finally, we remind that all loop-induced rates, branching ratios and total widths in our scan are evaluated with the spectrum tools described in Sec. 4.3.
4 Methodology and Constraints
In order to identify viable regions of parameter space within the GM model, we perform a numerical scan subject to both theoretical and experimental constraints. This section describes the methodology of the scan together with the various constraints imposed. The detailed numerical outcomes corresponding to these constraints will be presented in Sec. 5.
4.1 Experimental constraints
As discussed in Sec. 2, the GM model extends the SM by introducing several new scalar particles. Therefore, its predictions must be consistent with current experimental observations.
  • Electroweak precision measurements.  The effects of new scalar states in the GM model enter electroweak precision observables through their contributions to the self-energies of the $W$ and $Z$ bosons as well as neutral-current processes. At the one-loop level such oblique corrections are conveniently parametrized by the Peskin-Takeuchi variables $S$, $T$, and $U$ Peskin:1990zt ; Peskin:1991sw . The latest global fit based on precision electroweak data yields ParticleDataGroup:2024cfk
    $T=0.01\pm 0.12,\qquad S=-0.04\pm 0.10,\qquad U=-0.01\pm 0.09.$ (29)
    In addition, the $\rho$ parameter, which characterizes possible new sources of $\rm SU(2)$} breaking beyond the SM Higgs doublet, is constrained to be
    $\rho=1.00031\pm 0.00019.$ (30)
    Since custodial symmetry is preserved at tree level in the GM model due to its elegant scalar field structure, the slight deviation provides an intriguing hint about the GM model.
  • Higgs data fits.   The heavier singlet scalar $h$, identified as the SM-like Higgs boson at 125 GeV, must remain consistent with the combined ATLAS and CMS measurements at the $95\%$ confidence level. This requirement is implemented through package HiggsSignals-2.6.2 Bechtle:2013xfa ; Bechtle:2014ewa ; Bechtle:2020uwn ; Bahl:2022igd , demanding a $p$-value greater than $0.05$ and a total $\chi^{2}$ not exceeding the SM value by more than $\Delta\chi^{2}<6.18$ (corresponding to $95\%$ C.L. for two degrees of freedom).
    The couplings of the 125 GeV Higgs boson to fermions and vector bosons are also probed by global fits in the $\kappa$-framework. The ATLAS analysis reports
    $\begin{split}&\kappa_{tt}=0.94\pm 0.11,\quad\kappa_{bb}=0.89\pm 0.11,\quad\kappa_{\tau\tau}=0.93\pm 0.07,\\ &\kappa_{\mu\mu}=1.06^{+0.25}_{-0.30},\quad\kappa_{WW}=1.05\pm 0.06,\quad\kappa_{ZZ}=0.99\pm 0.06,\end{split}$ (31)
    while the CMS fit gives
    $\begin{split}&\kappa_{tt}=1.01\pm 0.11,\quad\kappa_{bb}=0.99\pm 0.16,\quad\kappa_{\tau\tau}=0.92\pm 0.08,\\ &\kappa_{\mu\mu}=1.12\pm 0.21,\quad\kappa_{WW}=1.02\pm 0.08,\quad\kappa_{ZZ}=1.04\pm 0.07.\end{split}$ (32)
    Combining the two sets of results, the approximate $1\sigma$ ranges adopted in the scan are
    $\begin{split}0.83\leq\kappa_{tt}\leq 1.12,\quad 0.78\leq\kappa_{bb}\leq 1.15,\quad 0.84\leq\kappa_{\tau\tau}\leq 1.00,\\ 0.81\leq\kappa_{\mu\mu}\leq 1.36,\quad 0.94\leq\kappa_{WW}\leq 1.11,\quad 0.93\leq\kappa_{ZZ}\leq 1.11.\end{split}$ (33)
    and these constraints are implemented in the scan through package HiggsSignals.
    In addition, the production and decay signal strengths of any extra Higgs states must not violate collider exclusion limits. These bounds, derived from LEP and LHC searches, are applied using HiggsBounds-5.10.0 Bahl:2022igd ; Bechtle:2008jh ; Bechtle:2011sb ; Bechtle:2013wla ; Bechtle:2020pkv , ensuring consistency with current experimental data.
  • $B$-physics observables. Rare decays of $B$ mesons provide sensitive probes of physics beyond the SM, since such processes are strongly suppressed within the SM itself. In particular, the branching ratios of $B_{s}\to\mu^{+}\mu^{-}$ and $B\to X_{s}\gamma$ are well measured and offer stringent constraints on extended Higgs sectors CMS:2014xfa . In our analysis we require
    $\begin{split}1.7\times 10^{-9}<&\mathcal{B}(B_{s}\to\mu^{+}\mu^{-})<4.5\times 10^{-9},\\ 2.99\times 10^{-4}<&\mathcal{B}(B\to X_{s}\gamma)<3.87\times 10^{-4},\end{split}$ (34)
    corresponding to the current $2\sigma$ experimental intervals.
4.2 Theoretical constraints
In addition to the above experimental limits, theoretical considerations further constrain the parameter space, ensuring the internal consistency of the GM model.
  • Perturbative unitarity.  Perturbative unitarity requires that the zeroth partial-wave amplitude $a_{0}$ of scalar-scalar scattering obey $|a_{0}|<1$, or equivalently ${\rm Re}(a_{0})<1/2$. In the GM model this translates into bounds on linear combinations of quartic couplings Hartling:2014zca ; Chowdhury:2024mfu ; Aoki:2007ah , leading to the conditions
    $\begin{split}\sqrt{(6\lambda_{1}-7\lambda_{3}-11\lambda_{4})^{2}+36\lambda_{2}^{2}}+\left|6\lambda_{1}+7\lambda_{3}+11\lambda_{4}\right|&<4\pi,\\ \sqrt{(2\lambda_{1}+\lambda_{3}-2\lambda_{4})^{2}+\lambda_{5}^{2}}+\left|2\lambda_{1}-\lambda_{3}+2\lambda_{4}\right|&<4\pi,\\ \left|2\lambda_{3}+\lambda_{4}\right|<\pi,\qquad\left|\lambda_{2}-\lambda_{5}\right|&<2\pi.\end{split}$ (35)
    These inequalities are applied to all scanned points to guarantee the perturbative consistency of the scalar sector.
  • Positive definiteness constraints.  It is well known that quantum corrections can significantly modify the shape of the scalar potential, potentially leading to violations of the tree-level bounded-from-below (BFB) conditions Hartling:2014zca . Such effects may induce deeper vacua at large field values or render the potential unbounded, implying that the electroweak vacuum preserving custodial symmetry may not be the true ground state. This situation is reminiscent of the metastability of the SM electroweak vacuum for a 125 GeV Higgs boson.
    To address this issue, we impose the so-called positive definiteness constraints Du:2024mry , which guarantee the stability of the custodial vacuum by requiring positivity of certain subdeterminants of the quartic coupling matrices in the one-loop RG-improved potential. We parameterize the one-loop RG-improved scalar potential by the dimensionless couplings $\lambda$, $\rho_{1,2,\dotsc,5}$ and $\sigma_{1,2,3,4}$ of the most general ${\rm SU(3)}_{C}\times{\rm SU(2)}_{L}\times{\rm U(1)}_{Y}$ invariant GM model. [^2] Dimensional parameters are eliminated via the minimisation conditions for the VEVs, and thus do not affect the large-field analysis relevant to positive definiteness. In the custodial-symmetric limit these couplings map onto the standard GM quartics $\lambda_{1,2,\dots,5}$ as
    $\begin{split}\lambda&=4\lambda_{1},\\ \rho_{1}&=4\lambda_{2}+6\lambda_{3},\quad\rho_{2}=-4\lambda_{3},\quad\rho_{3}=2(\lambda_{2}+\lambda_{3}),\quad\rho_{4}=4\lambda_{2},\quad\rho_{5}=4\lambda_{3},\\ \sigma_{1}&=4\lambda_{4}-\lambda_{5},\quad\sigma_{2}=2\lambda_{5},\quad\sigma_{3}=2\lambda_{4},\quad\sigma_{4}=\sqrt{2}\lambda_{5}.\end{split}$ (36)
    With these definitions and when the $\rho_{5}$ and $\sigma_{4}$ are small, the positive definiteness constraints then reduce to requiring that all leading principal minors of $\mathcal{V}$ are strictly positive,
    $\det\mathcal{V}_{n}>0\qquad(n=1,2,\dotsc,5),$ (37)
    where the $\mathcal{V}$ defined as
    $\mathcal{V}=\begin{pmatrix}\frac{1}{4}\lambda&\frac{1}{8}(2\sigma_{1}+\sigma_{2})&\frac{1}{4}\sigma_{1}&\frac{1}{2}\sigma_{3}~&\frac{1}{4}\sigma_{3}\\[4.0pt] \frac{1}{8}(2\sigma_{1}+\sigma_{2})&\frac{1}{2}(2\rho_{1}+\rho_{2})&\frac{3}{2}\rho_{2}&\rho_{4}&\frac{1}{2}\rho_{4}\\[4.0pt] \frac{1}{4}\sigma_{1}&\frac{3}{2}\rho_{2}&\rho_{1}+\rho_{2}&\rho_{4}&\frac{1}{2}\rho_{4}\\[4.0pt] \frac{1}{4}\sigma_{3}&\rho_{4}&\rho_{4}&2\rho_{3}&\rho_{3}\\[4.0pt] \frac{1}{2}\sigma_{3}&\frac{1}{2}\rho_{4}&\frac{1}{2}\rho_{4}&\rho_{3}&\frac{1}{2}\rho_{3}\end{pmatrix}$, (38)
    and $\mathcal{V}_{n}$ denotes the $n\times n$ upper-left submatrix of $\mathcal{V}$.
    It should be noted that the coupling parameters (such as $\lambda$, $\rho_{1,2,\dotsc,5}$, and $\sigma_{1,2,3,4}$, etc.) must be replaced by their renormalization group evolved values in regions of large field values. These evolved parameters can then be used to determine the vacuum structure of the GM model.
    We utilise the iterative method described in Du:2024mry to assess whether the parameter space that adequately explains the observed excess signals satisfies the positive definiteness constraints. The gauge coupling values at the EW scale employed in this study are derived from Antusch:2013jca . To ensure that the vacuum maintains custodial symmetry, the boundary conditions for the scalar quartic couplings at the EW scale, which are crucial for the RGEs, are selected in accordance with Eq. (36). To determine the scale dependence of the quartic couplings within the one-loop RG-improved tree-level scalar potential, it is essential to evolve the input parameters to an arbitrary scale, especially in regions of large field values, using the appropriate beta functions. In our detailed numerical analysis, to establish the positive definiteness constraints, we evolve the relevant quartic couplings and mass dimension parameters from the EW scale ($M_{Z}=91.2~{\rm GeV}$) up to the Planck scale ($M_{\rm Planck}\approx 1.12\times{10}^{19}~{\rm GeV}$) or to the lowest Landau pole scale, where any of the quartic couplings diverge.
4.3 Scanning setup
In order to realise the light custodial singlet scalar $H$ lie near $95.4~{\rm GeV}$ capable of accomodating the LHC hints in the $\gamma\gamma$ and $\tau\tau$ channels. In practice, we ultilise the package GMCALC Hartling:2014xma , a specifical GM model calculator. Alternatively, we use the SPheno package Porod:2003um ; Porod:2011nf in combination with SARAH Staub:2013tta ; Staub:2012pb ; Staub:2010jh ; Staub:2009bi ; Staub:2008uz to compute the GM model spectrum. We use the package HiggsSignals-2.6.2 Bechtle:2013xfa ; Bechtle:2014ewa ; Bechtle:2020uwn ; Bahl:2022igd to examine whether the properties of the next-to-lightest Higgs boson $h$ resemble those of the SM-like Higgs boson. The additional light Higgs boson $H$ is tested against collider constraints using the package HiggsBounds-5.10.0 Bahl:2022igd ; Bechtle:2008jh ; Bechtle:2011sb ; Bechtle:2013wla ; Bechtle:2020pkv .
We conduct a comprehensive scan of the input parameters in Eq. (20), applying all constraints described in Sec. 4.1 and Sec. 4.2, over the following ranges:
$\begin{split}-\sqrt{4\pi}&<{\lambda_{2,3,4,5}}<\sqrt{4\pi},\\ 50~\mathrm{GeV}&<m_{h,H}^{\rm In}<200~\mathrm{GeV},\\ 0<\sin{\theta}&<0.45,\quad 0<\alpha<\pi.\end{split}$ (39)
The two singlet Higgs states are restricted to
$94.4~{\rm GeV}<m_{H}<96.4~{\rm GeV},\quad 124.1~{\rm GeV}<m_{h}<126.1~{\rm GeV}.$ (40)
The $\chi^{2}$ function quantifying the ability of light custodial singlet scalar $H$ to explain the excesses in the $\gamma\gamma$ and $\tau\tau$ channels is given by
$\chi^{2}_{95}=\left(\frac{\mu_{\gamma\gamma}-0.24}{0.09}\right)^{2}+\left(\frac{\mu_{\tau\tau}-1.2}{0.5}\right)^{2}.$ (41)
In the absence of a published combined likelihood, correlations between $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ are neglected. Unless stated otherwise, points with $\chi^{2}_{95}<6.18$ (two observables) are regarded as consistent at the $2\sigma$ level.
5 Numerical results
Figure 1: Scan samples under successive constraints in the ($\mu_{\gamma\gamma}$, $\mu_{\tau\tau}$) panel with color coded by $\chi^{2}_{\rm 95}$. The left panel applies perturbative unitarity, the positive-definiteness vacuum-stability conditions, and passes the HiggsBounds/HiggsSignals tests. The right panel additionally imposes constraints from $B$-physics, the oblique parameters $(S,T,U)$, and the Higgs $\kappa$ fits.
Based on the scanned samples, the single light custodial singlet $H$ in the GM model near 95.4 GeV can account for the LHC $\gamma\gamma$ and $\tau\tau$ excesses at the $2\sigma$ level. Quantitatively, the scan covers the reported $2\sigma$ region in $\mu_{\gamma\gamma}$, while $\mu_{\tau\tau}$ typically peaks near 0.55 after all cuts. This underscores the lingering tension discussed at the end of Sec.  3 and leads to a definitive result.
Figure 1 shows the scan in the $(\mu_{\gamma\gamma},\mu_{\tau\tau})$ plane, with the color indicating $\chi^{2}_{95}$ defined in Eq. (41). The left panel applies perturbative unitarity, the positive-definiteness vacuum-stability conditions, and the HiggsBounds/HiggsSignals constraints. The right panel further incorporates the $B$-physics observations, electroweak precision constraints $(S,T,U)$, and fits to the SM-like Higgs $\kappa$ parameters. In the left panel, the samples are distributed almost uniformly across $\mu_{\gamma\gamma}\in[0.08,0.40]$, while $\mu_{\tau\tau}$ shows a different pattern. The lowest $\chi^{2}_{95}$ values are concentrated along a ridge where moderate values of $\mu_{\gamma\gamma}$ (approximately 0.20 to 0.3) coincide with $\mu_{\tau\tau}$ values around 0.50 to 0.54. The maximum value of $\mu_{\tau\tau}$ can reach up to 0.61. After incorporating $B$ physics, $(S,T,U)$, and the $\kappa$-constraints (right panel), the distribution becomes more compact: the high-$\mu_{\tau\tau}$ tail above approximately 0.55 and the extreme edges of $\mu_{\gamma\gamma}$ are eliminated. This behavior reflects the interplay between the charged-scalar loop enhancement of $H\to\gamma\gamma$ and the rescaling of fermionic couplings, which constrains $\mu_{\tau\tau}$ as the total width increases, as we discussed in Sec. 3.
Figure 2: Similar to Fig. 1: Top panel shows the ($\kappa_{ff}$, $\kappa_{VV}$) plane, with color indicating the $p$-value as calculated by HiggsSignals. The middle panel displays the ($\alpha$, $\mathcal{B}(B\to X_{s}\gamma)$) plane, with color representing $\mathcal{B}(B_{s}^{0}\to\mu^{+}\mu^{-})$. The bottom panel presents the distributions of the oblique parameters $S$, $T$, and $U$.
In Fig. 2, we present the distributions of the effective Higgs couplings $\kappa_{ff}^{h}$ and $\kappa_{VV}^{h}$ for the SM-like Higgs state $h$, the singlet mixing angle $\alpha$, and the oblique parameters $S$, $T$, and $U$ to illustrate their effects. Before applying these constraints, it is evident that the samples distributed into three distinct clusters. All three clusters require significant mixing between the doublet state $\phi$ and the singlet from the $\Delta$ field. Consequently, the mixing angle $\alpha$ must be approximately $\frac{\pi}{3}$ or $\frac{2\pi}{3}$, both of which correspond to similar values of $\left|\kappa_{ff}^{H}\right|^{2}$. In this case, the lighter $CP$-even Higgs state $H$ can acquire a sufficiently large Yukawa coupling, which in turn significantly affects its $ggF$ production rate and total decay width.
The first type of sample is characterised by $\alpha\approx 1.05$ ($\approx\frac{\pi}{3}$). For this type samples, although $\Gamma(H\to\tau\tau)$ is relatively large, the total width $\Gamma^{\rm tot}_{H}$ does not increase significantly for two reasons. First, with $m_{H}$ at only about 95 GeV, many decay channels—such as $t\bar{t}$, $WW$, and ZZZZ—are either kinematically forbidden or highly suppressed. Second, when $\alpha\approx 1$, the coupling of $H$ to gauge bosons is not particularly strong, so the total width does not increase substantially due to $VV$ modes. As a result, the denominator $\Gamma^{\rm tot}_{H}$ does not become large enough to "dilute" the enhancement in the numerator. Consequently, $\mu_{\tau\tau}$ can remain around $\sim 0.6$, which is higher than in the case with $\alpha\approx 2.1$. In contrast, in the branch with $\alpha\approx 2.1$, the gauge boson couplings of $H$ are closer to the SM structure. As a result, the contributions from $H\to VV$ and other lepton/quark channels to the total width are larger, leading to an increase in $\Gamma^{\rm tot}_{H}$ and consequently a suppression of $\mu_{\tau\tau}$. This highlights the tension between the SM-like Higgs state fit and the interpretation of the 95 GeV signal within the GM model. The first type of samples also predict a lower value of $\kappa^{h}_{VV}$ and are therefore completely excluded by Higgs effective coupling measurements, as shown in the top panel of Fig. 2.
Figure 3: Similar to Fig. 1: Top panel shows the ($m_{H},v_{\Delta}$) plane, with color indicating the $p_{\text{value}}$; bottom panel shows the ($m_{h},m_{H_{5}}$) plane, with color representing $m_{H_{3}}$.
To further illustrate the remaining two classes with $\alpha\sim 2.1$, we present plots of the samples in the ($m_{H}$, $v_{\Delta}$) plane, color-coded by $p$-value, and in the ($m_{h}$, $m_{H_{5}}$) plane, with colors representing $m_{H_{3}}$, as shown in Fig. 3. The second type of samples is characterized by $v_{\Delta}$ greater than 10 GeV, while the remaining samples have $v_{\Delta}$ less than 10 GeV. In the GM model, for the 125 GeV Higgs to exhibit SM-like behavior, its effective couplings to gauge bosons ($W$, $Z$) and fermions ($t$, $b$, $\tau$) must closely match their SM values. This requirement is directly linked to the doublet-triplet mixing angle $\theta$ and the triplet vacuum expectation value (VEV) $v_{\Delta}$. When $v_{\Delta}$ is small ($\lesssim 3$ GeV), the triplet component in the SM-like Higgs state $h$ is negligible, resulting in a significantly reduced coupling of the 125 GeV Higgs to $W$ and $Z$ bosons ($\kappa_{VV}^{h}<1$). This deviation undermines the SM-like behavior of the Higgs and creates tension with global $\kappa$ fits from the LHC. Conversely, when $v_{\Delta}$ is large ($\gtrsim 10$ GeV), the gauge boson coupling of the 125 GeV Higgs (see Eq. (22)) can be tuned to be very close to 1. In this regime, the decay and production rates of $h$ align with SM predictions, satisfying the constraints from HiggsSignals. In other words, a larger $v_{\Delta}$ (>10>10 GeV) can effectively "absorb" the effects of new physics, making the 125 GeV Higgs $h$ more SM-like. This trend is evident in numerical scans: when $v_{\Delta}$ falls within the [10, 15] GeV range, the surviving parameter points yield higher $p$-values (as reported by HiggsSignals), indicating better agreement with experimental data.
From the bottom panels of Fig. 2, The distributions of the oblique Peskin-Takeuchi parameters $T$, $S$, and $U$ differ significantly among the three sample classes. This is fairly straightforward. The $T$ parameter is often regarded as a sensitive probe of custodial symmetry breaking. In the GM model, custodial symmetry is preserved at tree level, so the primary contributions to $T$ come from mass splittings within the triplet and quintuplet states. For all samples, the custodial symmetry is nearly exact, resulting in $T\approx 0$. For larger $v_{\Delta}$ or varying $\alpha$, mixing between the light scalar $H$ and the 125 GeV Higgs can shift the masses of certain scalars up or down, leading to mass differences within custodial multiplets and thus deviations in $T$. The $S$ parameter represents corrections to the $Z$ boson self-energy and its mixing, arising from new physics states. In GM model, $S$ parameter is sensitive to light charged scalars ($H_{3}^{\pm}$, $H_{5}^{\pm}$, $H_{5}^{\pm\pm}$), which shown in Fig 3. When the light scalar $H$ has a larger triplet component, for example in the region $\alpha\approx 1$, loop effects are enhanced, leading to more significant deviations in $S$. Conversely, if the 125 GeV Higgs remains SM-like, for example $\alpha\simeq 2.1$ and large $v_{\Delta}$, the new states are heavier or have smaller mixing, resulting in smaller corrections to $S$.
Let us turn to the $B$-rare decays. The $B$-rare decay processes are highly suppressed in the SM, dominated mainly by top quark and $W$ boson loops. While in the GM model, new charged scalars such as $H_{3}^{\pm}$, $H_{5}^{\pm}$, and $H_{5}^{\pm\pm}$ are introduced, which can also participate in these loop diagrams. The masses of these charged scalars and their couplings to fermions (which depend on $\alpha$ and $v_{\Delta}$) vary across different parameter regions, leading to different corrections to the $B$-rare decay amplitudes. In the so-called "alignment limit" of the GM model, the constraints from rare B-meson decays are not very stringent. As shown in these two panels, all sample points fall within the allowed range. Therefore, as discussed earlier, although the branching ratios for rare $B$ decays are of similar magnitude, they can still be classified into three distinct groups.
Finally, the large signal strength observed in the di-tau excess appears somewhat inconsistent with recent experimental constraints from searches for Higgs boson production in association with either $t\bar{t}$ or a $Z$ boson, followed by decay into tau pairs, as reported by CMS CMS:2024ulc . This tension is further heightened by LEP results from searches for the process $e^{+}e^{-}\to Z(\phi\to\tau^{+}\tau^{-})$ LEP:2003ing . Although the observed value is still compatible with current experimental limits, it will be thoroughly tested at the HL-LHC ZurbanoFernandez:2020cco ; LHCb:2018roe (with an expected precision of $3.4\%$ for $\kappa_{t}$) and at CEPC CEPCPhysicsStudyGroup:2022uwl ; CEPCStudyGroup:2023quu ; Ai:2024nmn ; Antusch:2025lpm (with an expected precision of less than $1\%$ for $\kappa_{t}$). Recent global fits of GM models further support this conclusion, indicating that the realistic signal strength in the di-tau channel should be significantly lower. It is also worth noting that the ATLAS collaboration has not yet published results from searches in the di-tau final state for masses around 95 GeV. Forthcoming results from ATLAS are therefore expected to clarify the situation, potentially reducing the central value of the $\tau\tau$ excess and providing a clearer understanding of the underlying physics.
6 Conclusions
In this work, we showed that the GM model can provide a unified interpretation of the LHC hints near 95 GeV in the $\gamma\gamma$ and $\tau\tau$ channels with a single light $CP$-even custodial singlet. Implementing the RG-improved positive-definiteness conditions for vacuum stability enlarges the viable parameter space relative to traditional bounded-from-below requirements, while remaining consistent with perturbative unitarity, electroweak precision data, $B$-physics observables, direct searches, and Higgs signal-strength measurements.
The surviving parameter region displays distinctive features that make the scenario predictive: singlet mixing is restricted to narrow intervals; the triplet VEV $v_{\Delta}$ is at the few 10 GeV scale; and the custodial multiplets exhibit a mass hierarchy with $m_{H_{3}}<m_{H_{5}}$ in the ${\cal O}(10^{2})~{\rm GeV}$ order. The $\gamma\gamma$ production rate is enhanced by charged-scalar loops, most notably from the doubly charged state $H_{5}^{\pm\pm}$, which interfere constructively with the SM contributions. In contrast, the $\tau\tau$ signal strength is determined by the universal rescaling of fermionic couplings and the total decay width. Consequently, the scan readily reproduces the reported $2\sigma$ interval for $\mu_{\gamma\gamma}$, while $\mu_{\tau\tau}$ typically remains below the experimental central value but still within approximately $2\sigma$ compatibility.
These features make the framework highly testable in future experiments. At the HL-LHC, updated searches in the $\gamma\gamma$ and $\tau\tau$ final states, along with direct probes of $H_{5}^{\pm\pm}$, $H_{5}^{\pm}$, and $H_{3}^{\pm}$, will be able to explore much of the remaining parameter space. Future lepton colliders, such as CEPC and FCC-ee, will further offer decisive tests of this scenario.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grant Nos. 12447167 and 12275067), by the Joint Fund of Henan Province Science and Technology R&\&D Program (Grant Nos.225200810030, 225200810092 and 225200810014), by the Startup Research Fund of Henan Academy of Sciences (Grant No. 231820011), by the Basic Research Fund of Henan Academy of Sciences (Grant No. 240620006), by the Graduate Innovation Fund of Henan Academy of Sciences (Grant No. 243320031), by the Science and Technology Innovation Leading Talent Support Program of Henan Province, by the National Key R&\&D Program of China (Grant No. 2023YFA1606000) and by the Science and Technology Innovation Leading Talent Support Program of Henan Province (Grant No. 254000510039). PZ is supported by the ARC Discovery Project DP220100007 and by the Centre for the Subatomic Structure of Matter (CSSM).