### v2

# Unified interpretation of 95 GeV di-photon and di-tau Excesses in the Georgi-Machacek Model

Abstract
We revisit the 95 GeV excesses in the $\gamma\gamma$ and $\tau\tau$ channels in the Georgi–Machacek model, where a single light $CP$-even custodial singlet $H$ can account for both hints. Using one-loop renormalization group-improved effective potential and positive-definiteness conditions for vacuum stability, together with perturbative unitarity, electroweak precision tests, $B$-physics, and Higgs data, we identify a narrow but viable parameter region. In this region, charged and doubly charged scalars enhance $H\to\gamma\gamma$, while the fermionic rescaling controls $H\to\tau\tau$, yielding combined signal strengths compatible at the $2\sigma$ level. The scenario predicts characteristic patterns in the singlet mixing and triplet vacuum expectation values and is highly testable at the HL-LHC and future lepton colliders via precision $\kappa_{V}$ measurements and direct exotic searches.
1 Introduction
The discovery of the 125 GeV Higgs boson by the ATLAS ATLAS:2012ae and CMS CMS:2012zhx collaborations in 2012 confirms the Standard Model (SM) framework of electroweak symmetry breaking (EWSB). Nevertheless, theoretical motivations and the sizable uncertainties in current Higgs data fits still leave ample room for an extended Higgs sector ATLAS:2019nkf . Consequently, the search for additional scalar fields remains a crucial objective.
[…]
The Georgi-Machacek (GM) model Georgi:1985nv ; Chanowitz:1985ug ; Gunion:1989ci ; Gunion:1990dt ; Gunion:1989we ; Haber:1999zh ; Godfrey:2010qb ; Logan:2010en ; Low:2012rj ; Killick:2013mya ; Efrati:2014uta ; Degrande:2015xnm extends the SM by introducing a complex ${\rm SU(2)}_{L}$ triplet scalar $\chi$ with hypercharge +1+1 and a real ${\rm SU(2)}_{L}$ triplet scalar $\xi$ with hypercharge 0. The field content is arranged such that custodial symmetry is preserved at tree level after EWSB, provided the triplet vacuum expectation values (VEVs) are aligned, $v_{\chi}=v_{\xi}\equiv v_{\Delta}$. This alignment permits $v_{\Delta}$ values up to several tens GeV, significantly larger than in generic triplet extensions, while maintaining $\rho_{\rm tree}=1$. The model predicts nine additional scalar states, including a characteristic doubly charged Higgs boson, resulting in a rich phenomenology and distinctive collider signatures Chiang:2012cn ; Kanemura:2013mc ; Englert:2013zpa ; Englert:2013wga ; Godunov:2014waa ; Chiang:2015kka ; Chiang:2015amq ; Yu:2016skv ; Cao:2016rht ; Chang:2017niy ; Zhang:2017och ; Li:2017daq ; Sun:2017mue ; Azuelos:2017dqw ; Cao:2018xwy ; Yang:2018pzt ; Ghosh:2019qie ; Ismail:2020zoz ; Zhu:2020bew ; Wang:2022okq ; Ghosh:2022wbe ; deLima:2022yvn ; Ahriche:2022aoj ; Ghosh:2022bvz ; Chakraborti:2023mya ; Ghosh:2023izq ; Bhattacharya:2024gyh ; Lu:2024ade ; Ghosh:2025htt ; Ashanujjaman:2025puu ; Ghosh:2025hyy ; Chiang:2025lab ; Chiang:2018cgb . Notably, the GM model allows the couplings of the SM-like Higgs boson to weak gauge bosons to exceed their SM values Falkowski:2012vh ; Chang:2012gn ; Chiang:2013rua .
The extended scalar sector improves vacuum stability and can induce strong first-order electroweak phase transitions (EWPT), potentially generating gravitational wave signals accessible to future experiments Chiang:2014hia ; Chen:2022zsh ; Zhou:2018zli ; Bian:2019bsn ; Lu:2025vif . If lepton number violation is present, triplet-lepton interactions can generate neutrino masses Chen:2020ark ; Giarnetti:2023dcr ; Giarnetti:2023osf ; Du:2022brr . The GM framework has also been employed to address the $W$-mass anomaly reported by CDF-II Du:2022brr ; Mondal:2022xdy ; Chen:2022ocr ; Ghosh:2022zqs , especially when custodial symmetry breaking effects are included Blasi:2017xmc ; Keeshan:2018ypw ; Song:2022jns . Furthermore, the light custodial singlet scalar of the GM model constitutes a natural candidate to interpret the excesses observed near 95 GeV Ahriche:2023wkj ; Chen:2023bqr ; Du:2025eop ; Ghosh:2025ebc ; Mondal:2025tzi . A unified explanation for the excesses observed in the $\gamma\gamma$, $\tau\tau$, and $b\bar{b}$ channels requires a near-degenerate CPCP-even/CPCP-odd ‘twin-peak’ structure Ahriche:2023wkj ; with only a single light $CP$-even state, the $\tau\tau$ excess is in tension with the $\gamma\gamma$ and $b\bar{b}$ channels Ahriche:2023wkj ; Chen:2023bqr ; Du:2025eop . Recent analyses, however, strongly disfavor the LEP $b\bar{b}$ excess as evidence for a new scalar Janot:2024ryq , motivating a focus on the LHC-driven $\gamma\gamma$ and $\tau\tau$ channels.
The positive-definiteness vacuum stability conditions for the GM model were proposed in Du:2024mry , which preclude the existence of deeper vacua in large field value regions. Moving beyond the conventional tree-level analysis, the study employs the one-loop renormalization group-improved (RG-improved) effective potential, along with a novel criterion for the positive definiteness of homogeneous polynomials in multiple variables—a refinement necessitated by loop-induced custodial symmetry breaking effects. Through numerical analysis, it is shown that the quartic coupling ranges permitted by the positive-definiteness requirements significantly deviate from those derived from tree-level bounded-from-below conditions. Notably, certain parameter regions previously excluded by tree-level criteria become viable when the one-loop RG-improved scalar potential is considered. Conversely, some regions of parameter space that satisfy tree-level bounded-from-below constraints at the electroweak scale are ultimately ruled out by the positive-definiteness requirements of the effective potential in large field value regimes.
In this work, the newly proposed positive-definiteness vacuum stability conditions based on the one-loop RG-improved effective potential expand the viable parameter space beyond that allowed by traditional bounded-from-below requirements. This enables a unified interpretation of both excesses within a single $CP$-even Higgs-boson scenario, and we therefore focus on the unified interpretation of $\gamma\gamma$ and $\tau\tau$ resonance signals within the GM model.
This work is organized as follows. In Sec. 2, we provide a brief review of the GM model. A short discussion of the di-photon and di-tau excess signals around 95 GeV is presented in Sec. 3. Theoretical and experimental constraints relevant to realistic studies within the GM model are detailed in Sec. 4. Numerical results are displayed in Sec. 5. Finally, we present our conclusions in Sec. 6.
2 The Georgi-Machacek model
The GM model extends the SM scalar sector by introducing additional triplet fields. It contains the familiar ${\rm SU(2)}_{L}$ doublet $({\phi}^{+},~{\phi}^{0})$ with hypercharge $1/2$, a complex triplet $({\chi}^{++},~{\chi}^{+},~{\chi}^{0})$ with hypercharge +1+1, and a real triplet $({\xi}^{+},~{\xi}^{0},~-{\xi}^{+*})$ with zero hypercharge. This field structure enables custodial ${\rm SU(2)}_{V}$ symmetry at tree level and leads to distinctive scalar phenomenology.
To clarify custodial symmetry, the doublet and triplets are arranged in matrices $\Phi$ and $\Delta$, respectively:
$\Phi=\begin{pmatrix}\phi^{0*}&\phi^{+}\\ \phi^{-}&\phi^{0}\end{pmatrix},\quad\Delta=\begin{pmatrix}\chi^{0*}&\xi^{+}&\chi^{++}\\ \chi^{-}&\xi^{0}&\chi^{+}\\ \chi^{--}&\xi^{-}&\chi^{0}\end{pmatrix}$. (1)
Then the most general scalar potential for the GM scalar sector, constructed to be invariant under a global ${\rm SU(2)}_{L}\times{\rm SU(2)}_{R}$ symmetry, hence also under its gauged subgroup ${\rm SU(2)}_{L}\times{\rm U(1)}_{Y}$, reads
$\begin{split}V(\Phi,\Delta)=&\frac{1}{2}m_{\Phi}^{2}{\rm tr}\left[\Phi^{\dagger}\Phi\right]+\frac{1}{2}m_{\Delta}^{2}{\rm tr}\left[\Delta^{\dagger}\Delta\right]\\ &+\lambda_{1}\left({\rm tr}\left[\Phi^{\dagger}\Phi\right]\right)^{2}+\lambda_{2}\left({\rm tr}\left[\Delta^{\dagger}\Delta\right]\right)^{2}+\lambda_{3}{\rm tr}\left[\left(\Delta^{\dagger}\Delta\right)^{2}\right]\\ &+\lambda_{4}{\rm tr}\left[\Phi^{\dagger}\Phi\right]{\rm tr}\left[\Delta^{\dagger}\Delta\right]+\lambda_{5}{\rm tr}\left[\Phi^{\dagger}\frac{\sigma^{a}}{2}\Phi\frac{\sigma^{b}}{2}\right]{\rm tr}\left[\Delta^{\dagger}T^{a}\Delta T^{b}\right]\\ &+\mu_{1}{\rm tr}\left[\Phi^{\dagger}\frac{\sigma^{a}}{2}\Phi\frac{\sigma^{b}}{2}\right](P^{\dagger}\Delta P)_{ab}+\mu_{2}{\rm tr}\left[\Delta^{\dagger}T^{a}\Delta T^{b}\right](P^{\dagger}\Delta P)_{ab},\end{split}$ (2)
Here, $\sigma^{i}$ are Pauli matrices, and $T^{i}$ are the $\rm SU(2)$} generators in the triplet representation:
$T^{1}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&1&0\\ 1&0&1\\ 0&1&0\end{pmatrix},\quad T^{2}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&-\mathrm{i}&0\\ \mathrm{i}&0&-\mathrm{i}\\ 0&\mathrm{i}&0\end{pmatrix},\quad T^{3}=\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&-1\end{pmatrix}$. (3)
The $P$ matrix rotates the triplet fields $\Delta$ to the Cartesian basis
$P=$ $\frac{1}{\sqrt{2}}\begin{pmatrix}-1&\mathrm{i}&0\\ 0&0&\sqrt{2}\\ 1&\mathrm{i}&0\end{pmatrix}$. (4)
It should be noted that the last two terms in Eq. (2) vanish under a $\mathbb{Z}_{2}$ symmetry. However, we retain them in this work.
Under ${\rm SU(2)}_{L}\times{\rm SU(2)}_{R}$, the fields $\Phi$ and $\Delta$ transform as
$\Phi\to U_{L}\Phi U_{R}^{\dagger},\quad\Delta\to U_{L}\Delta U_{R}^{\dagger}.$ (5)
Here, $U_{L,R}$ denotes $\exp(\mathrm{i}\theta^{a}_{L,R}\sigma^{a})$ for the doublet and $\exp(\mathrm{i}\theta^{a}_{L,R}T^{a})$ for the triplet, respectively. The SM electroweak gauge group ${\rm SU(2)}_{L}\times{\rm U(1)}_{Y}$ is embedded in the global ${\rm SU(2)}_{L}\times{\rm SU(2)}_{R}$ symmetry, with ${\rm SU(2)}_{L}$ identified by the generators $T_{L}^{i}$ and the hypercharge ${\rm U(1)}_{Y}$ associated with the generator $T_{R}^{3}$.
After the EWSB, the neutral scalar fields can be parameterized as
$\begin{split}{\xi}^{0}&\to v_{\xi}+h_{\xi},\\ {\phi}^{0}&\to\frac{1}{\sqrt{2}}\left(v_{\phi}+h_{\phi}+ia_{\phi}\right)\\ {\chi}^{0}&\to v_{\chi}+\frac{1}{\sqrt{2}}\left(h_{\chi}+ia_{\chi}\right).\end{split}$ (6)
Here, the parameters $v_{\xi}$, $v_{\phi}$ and $v_{\chi}$ represent the VEVs of the corresponding neutral components of these scalar fields. Then the tree level $\rho$ parameter of the GM model can be expressed as
$\rho_{\rm tree}\equiv\frac{m^{2}_{W}}{m_{Z}^{2}{\cos^{2}{\theta_{W}}}}=\frac{v_{\phi}^{2}+4v_{\chi}^{2}+4v_{\xi}^{2}}{v_{\phi}^{2}+8v_{\chi}^{2}}.$ (7)
Therefore, when the triplet VEVs are aligned, $v_{\chi}=v_{\xi}\equiv v_{\Delta}$, the vacuum remains invariant only under the diagonal (custodial) subgroup ${\rm SU(2)}_{V}\subset{\rm SU(2)}_{L}\times{\rm SU(2)}_{R}$ (i.e., when $U_{L}=U_{R}$). As a result, the tree-level $\rho$ parameter automatically satisfies $\rho_{\rm tree}=1$. In this alignment limit, the VEVs of the scalar fields take the form
$\langle\Phi\rangle=\frac{v_{\phi}}{\sqrt{2}}{\bf I}_{2\times 2},\quad\text{and}\quad\langle\Delta\rangle=v_{\Delta}{\bf I}_{3\times 3}.$ (8)
The condition for EWSB is given by
$v^{2}=v_{\phi}^{2}+8v_{\Delta}^{2}\simeq\left(246~{\rm GeV}\right)^{2}.$ (9)
The physical scalars are then decomposed into irreducible representations according to the custodial ${\rm SU(2)}_{V}$ symmetry. The doublet $\Phi$ (interpreted as a ${\bf 2}\otimes{\bf 2}$ representation) decomposes into ${\bf 1}\oplus{\bf 3}$, while the triplet $\Delta$ (as a ${\bf 3}\otimes{\bf 3}$ representation) decomposes into ${\bf 1}\oplus{\bf 3}\oplus{\bf 5}$. One triplet, $G_{3}$, provides the Nambu-Goldstone bosons eaten by $W$ and $Z$ bosons
$G_{3}^{\pm}=\cos{\theta}\phi^{\pm}+\frac{\sin{\theta}}{\sqrt{2}}\left(\chi^{\pm}+\xi^{\pm}\right),\quad G_{3}^{0}=\cos{\theta}{a_{\phi}}+\sin{\theta}a_{\chi},$ (10)
where the mixing parameter $\theta$ is defined by
$\tan{\theta}\equiv 2\sqrt{2}v_{\Delta}/v_{\phi}.$ (11)
The singlet $H_{1}^{0}$ (originating from $\Delta$ field), the triplet ($H_{3}^{\pm}$ and $H_{3}^{0}$), and the quintuplet ($H_{5}^{\pm\pm}$, $H_{5}^{\pm}$ and $H_{5}^{0}$) are given by
$\begin{split}H^{0}_{1}&=\sqrt{\frac{2}{3}}h_{\chi}+\sqrt{\frac{1}{3}}h_{\xi},\\ H_{3}^{\pm}&=-\sin{\theta}\phi^{\pm}+\frac{\cos{\theta}}{\sqrt{2}}\left(\chi^{\pm}+\xi^{\pm}\right),\quad H_{3}^{0}=\cos{\theta}a_{\chi}-\sin{\theta}a_{\phi},\\ H_{5}^{\pm\pm}&=\chi^{\pm\pm},\quad H_{5}^{\pm}=\frac{1}{\sqrt{2}}\left(\chi^{\pm}-\xi^{\pm}\right),\quad H_{5}^{0}=\sqrt{\frac{2}{3}}h_{\xi}-\sqrt{\frac{1}{3}}h_{\chi}.\end{split}$ (12)
In the basis of $(h_{\phi},H_{1}^{0},H_{5}^{0})$, the mass matrix of $CP$-even Higgs states is given by
$M^{2}_{CP\text{-even}}=\begin{pmatrix}M_{11}^{2}&M_{12}^{2}&0\\ M_{12}^{2}&M_{22}^{2}&0\\ 0&0&m_{H_{5}}^{2}\end{pmatrix}$, (13)
where the elements of the $2\times 2$ part are
$\begin{split}{M}_{11}^{2}&=8\lambda_{1}v_{\phi}^{2},\\ {M}_{12}^{2}={M}_{21}^{2}&=\frac{\sqrt{3}}{2}v_{\phi}\left(4\left(2\lambda_{2}+\lambda_{5}\right)v_{\Delta}+\mu_{1}\right),\\ {M}_{22}^{2}&=8\left(\lambda_{3}+4\lambda_{4}\right)v_{\Delta}^{2}-\frac{v_{\phi}^{2}}{4v_{\Delta}}\mu_{1}+6v_{\Delta}\mu_{2}.\end{split}$ (14)
The mass eigenstates $h$ and $H$ are defined by
$h=\cos{\alpha}h_{\phi}-\sin{\alpha}{H^{0}_{1}},\quad H=\sin{\alpha}h_{\phi}+\cos{\alpha}{H^{0}_{1}},$ (15)
where the mixing angle $\alpha$ is determined by
$\tan{2\alpha}=\frac{2M_{12}^{2}}{M_{22}^{2}-M_{11}^{2}},\quad\alpha\in\left(-\pi/2,\pi/2\right)$ (16)
The masses of all Higgs mass eigenstates can then be expressed as
$\begin{split}m_{h}^{2}&=M_{11}^{2}\cos^{2}{\alpha}+M_{22}^{2}\sin^{2}{\alpha}+M_{12}^{2}\sin{2\alpha},\\ m_{H}^{2}&=M_{11}^{2}\sin^{2}{\alpha}+M_{22}^{2}\cos^{2}{\alpha}-M_{12}^{2}\sin{2\alpha},\\ m_{H_{3}}^{2}\equiv m_{H_{3}^{\pm}}^{2}=m_{H_{3}^{0}}^{2}&=-\frac{v^{2}}{4v_{\Delta}}\mu_{1}-\frac{1}{2}\lambda_{5}v^{2},\\ m_{H_{5}}^{2}\equiv m_{H_{5}^{\pm\pm}}^{2}=m_{H_{5}^{\pm}}^{2}=m_{H_{5}^{0}}^{2}&=-\frac{v_{\phi}^{2}}{4v_{\Delta}}\mu_{1}-12v_{\Delta}\mu_{2}+8\lambda_{3}v_{\Delta}^{2}-\frac{3}{2}\lambda_{5}v_{\phi}^{2}.\end{split}$ (17)
By applying the tadpole conditions,
$\frac{\partial V}{\partial h_{\phi}}\bigg|_{0}=0,\quad\frac{\partial V}{\partial h_{\xi}}\bigg|_{0}=0,\quad\frac{\partial V}{\partial h_{\chi}}\bigg|_{0}=0,$ (18)
the mass parameters $m_{\phi}^{2}$ and $m_{\Delta}^{2}$ in Eq. (2) can be eliminated as follows:
$\begin{split}m_{\phi}^{2}&=-4\lambda_{1}v_{\phi}^{2}-\left(\frac{3}{2}\mu_{1}+6\lambda_{4}+3\lambda_{5}\right)v_{\Delta}^{2},\\ m_{\Delta}^{2}&=-(12\lambda_{2}+4\lambda_{3})v_{\Delta}^{2}-\left(\frac{1}{4v_{\Delta}}\mu_{1}+2\lambda_{4}+\lambda_{5}\right)v_{\phi}^{2}-6v_{\Delta}\mu_{2}.\end{split}$ (19)
In summary, the free parameters of the GM model consist of five dimensionless couplings $\lambda_{1,2,3,4,5}$, two dimensional parameters $\mu_{1,2}$, and one mixing parameter $\theta$. In practice, we adopt the following set of free parameters in this work:
$\lambda_{2,3,4,5},\quad m_{h}^{\rm In},\quad m_{H}^{\rm In},\quad\alpha,\quad\sin{\theta},$ (20)
where $\mu_{1,2}$ are substituted by $m_{h,H}^{\rm In}$, and $\lambda_{1}$ is traded for $\alpha$. In the following sections, $m_{h,H}$ without a superscript refers to the physical pole mass as calculated by the spectrum calculator.
3 95 GeV excesses in the GM Model
In the setup introduced in Sec. 2, the GM model contains two light singlet scalar fields. In this work, the heavier state, denoted as $h$, is identified with the $125~\mathrm{GeV}$ SM-like Higgs boson. The other state, labeled as $H$, serves as a candidate to explain the observed excesses around $95~\mathrm{GeV}$. The couplings of these Higgs states to SM fermions ($f$) and weak gauge bosons ($V=W,Z$) are modified as follows:
$\begin{split}g_{hff}&=\kappa_{F}\times g_{hff}^{\rm SM},\quad g_{hVV}=\kappa_{V}\times g_{hVV}^{\rm SM},\\ g_{Hff}&=\kappa_{F}^{H}\times g_{hff}^{\rm SM},\quad g_{HVV}=\kappa_{V}^{H}\times g_{hVV}^{\rm SM},\end{split}$ (21)
with the effective Higgs couplings $\kappa$ in the lowest order
[…]
Using this framework, the GM predictions for the di-photon signal strength normalized to the SM prediction $\mu_{\gamma\gamma}$ are given by
$\begin{split}\mu_{\gamma\gamma}&\equiv\frac{\sigma_{\rm GM}\left(pp\to H_{95}\to\gamma\gamma\right)}{\sigma_{\rm SM}\left(pp\to h_{95}\to\gamma\gamma\right)}\\ &\simeq\frac{\sigma_{\rm GM}\left(gg\to H_{95}\right)}{\sigma_{\rm SM}\left(gg\to h_{95}\right)}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to\gamma\gamma\right)}{\mathcal{B}_{\rm SM}\left(h_{95}\to\gamma\gamma\right)}\\ &=\left|\kappa_{F}^{H}\right|^{2}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to\gamma\gamma\right)}{1.39\times 10^{-3}},\end{split}$ (23)
where the mass of the Higgs boson is fixed around 95 GeV (denoted by $H_{95}$ and $h_{95}$), and the subscript $\rm GM$ ($\rm SM$) denotes the GM (SM) prediction for the inclusive production rate of $H_{95}$ ($h_{95}$). The specific value $1.39\times 10^{-3}$ is the SM Higgs branching ratio at 95 GeV calculated by LHC Higgs Cross Section Working Group LHCHiggsCrossSectionWorkingGroup:2013rie . The second line approximation arises because the production of the SM Higgs boson is dominated by the gluon-gluon fusion process, which accounts for approximately 89% of the total Higgs production at a mass of 95 GeV LHCHiggsCrossSectionWorkingGroup:2013rie . The third line arises because the gluon-gluon fusion process ($gg\to h$) is mediated by a colored-fermion loop. As a result, the effective coupling $\kappa^{H}_{gg}$ equals $\kappa^{H}_{F}$, since the fermionic couplings of $H$ scale universally as $\kappa_{F}^{H}$ in lowest order, as shown in Eq. (22). Similarly, the GM predictions for the signal strengths in the $\tau\tau$ and $bb$ channels are given by
$\begin{split}\mu_{bb}&\equiv\frac{\sigma_{\rm GM}\left(e^{+}e^{-}\to ZH_{95}(H_{95}\to bb)\right)}{\sigma_{\rm SM}\left(e^{+}e^{-}\to Zh_{95}(h_{95}\to bb)\right)}\\ &=\left|\kappa_{V}^{H}\right|^{2}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to bb\right)}{0.801},\\ \mu_{\tau\tau}&\equiv\frac{\sigma_{\rm GM}\left(pp\to H_{95}\to\tau\tau\right)}{\sigma_{\rm SM}\left(pp\to h_{95}\to\tau\tau\right)}\\ &=\left|\kappa_{F}^{H}\right|^{2}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to\tau\tau\right)}{8.32\times 10^{-2}}.\end{split}$ (24)
From the above definitions, all signal strengths are expressed in terms of the ratio $\sigma_{\rm GM}/\sigma_{\rm SM}$, so QCD corrections to the numerator and denominator largely cancel. This helps reduce the theoretical uncertainty in the predicted signal strengths.
Now, we scrutinize the branching ratios. The partial decay width of the lighter scalar $H$ into SM final states can be expressed as
$\Gamma(H\to XX)=|\kappa^{H}_{X}|^{2}\times\Gamma_{\rm SM}(h\to XX)\big|_{m_{h}=m_{H}},$ (25)
where $\Gamma_{\rm SM}(h\to XX)\big|_{m_{h}=m_{H}}$ is the SM Higgs partial decay width evaluated at mass $m_{H}$. Therefore, the total decay width of custodial singlet state $H$ can be written as
$\Gamma^{\rm tot}_{H}=\Gamma^{\rm SM}_{h}\sum_{X}\left|\kappa^{H}_{X}\right|^{2}\mathcal{B}_{\rm SM}\left(h\to XX\right),$ (26)
where the SM Higgs total decay width $\Gamma^{\rm SM}_{h}$ and branching ratios $\mathcal{B}_{\rm SM}\left(h\to XX\right)$ are evaluated at $m_{H}$. The couplings between the neutral scalars $H$ and $h$ and photons are induced by charged-particle loops. In the SM, Higgs-boson decay into two photons is mediated by $W$-boson and heavy charged-fermion (top-quark-dominated) loops. In the GM model, loops of heavy charged Higgs states $S_{i}$ (including $H_{5}^{\pm\pm}$, $H_{5}^{\pm}$, and $H_{3}^{\pm}$) can significantly enhance the partial decay width $\Gamma(H\to\gamma\gamma)$. The corresponding effective coupling $\kappa^{H}_{\gamma\gamma}$ can be expressed as
$\small\kappa^{H}_{\gamma}=\frac{\sum_{i}\dfrac{g_{HS_{i}S_{i}}}{2m_{S_{i}}^{2}}vQ_{S_{i}}^{2}A_{0}\left(\frac{m_{H}^{2}}{4m_{S_{i}}^{2}}\right)+\kappa_{V}^{H}A_{1}\left(\frac{m_{H}^{2}}{4m_{W}^{2}}\right)+\kappa_{F}^{H}~\dfrac{4}{3}A_{1/2}\left(\frac{m_{H}^{2}}{4m_{t}^{2}}\right)}{A_{1}\left(\frac{m_{H}^{2}}{4m_{W}^{2}}\right)+\dfrac{4}{3}A_{1/2}\left(\frac{m_{H}^{2}}{4m_{t}^{2}}\right)},$ (27)
where $g_{HS_{i}S_{i}}$ are the scalar trilinear couplings of $H$ to the charged states $S_{i}$, and the loop functions $A_{0,1/2,1}$ can be found in the literature Djouadi:2005gi ; Djouadi:2005gj ; Gunion:1989we . For an $H$ boson with mass $95~{\rm GeV}$, $A_{1}\simeq-7.6$ and $A_{1/2}\approx 4/3$. The factor $A_{0}$ reaches its maximum value of $1.5+\mathrm{i}$ at $m_{H}=2m_{S_{i}}$, while in the limit of very heavy loop charged scalars, $A_{0}(0)\simeq 1/3$.
Then the signal strength in Eq. (23) and Eq. (24) can be rewritten as:
$\begin{split}\mu_{\gamma\gamma}&=\left|\kappa^{H}_{F}\right|^{2}\left|\kappa^{H}_{\gamma}\right|^{2}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}}\\ \mu_{bb}&=\left|\kappa^{H}_{V}\right|^{2}\left|\kappa^{H}_{F}\right|^{2}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}},\\ \mu_{\tau\tau}&=\left|\kappa^{H}_{F}\right|^{4}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}}.\end{split}$ (28)
It is clear that all three signal strengths share a common factor, $\left|\kappa^{H}_{F}\right|^{2}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}}$. This introduces tension in attributing all observed excesses to a single source. Charged scalar states in the GM model can provide positive contributions to the $\gamma\gamma$ excess, especially the doubly charged $H_{5}^{\pm\pm}$ with $Q_{S}^{2}=4$. Ref. Chen:2023bqr investigates different possibilities for $H$ to explain $\gamma\gamma$, $\gamma\gamma+bb$, and $\gamma\gamma+bb+\tau\tau$. However, recent dedicated reanalyses argue that this feature is unlikely to originate from a genuine new scalar resonance Janot:2024ryq . Given the inferior di-jet mass resolution at LEP and the updated conclusions, we do not attempt to accommodate the $b\bar{b}$ hint in this work. Instead, we focus on the two LHC-driven channels ($\gamma\gamma$ and $\tau\tau$) where the experimental systematics and resolutions are better controlled in the relevant mass window. Finally, we note that all loop-induced rates, branching ratios and total widths in our scan are evaluated with the spectrum tools described in Sec. 4.
4 Methodology and Constraints
In order to realize a light custodial singlet scalar $H$ near $95.4~{\rm GeV}$ that is capable of accommodating the LHC hints in the $\gamma\gamma$ and $\tau\tau$ channels, we perform a comprehensive parameter scan. This section describes the methodology of the scan together with the various constraints imposed. The detailed numerical outcomes corresponding to these constraints will be presented in Sec. 5.
In practice, we utilize the package GMCALC Hartling:2014xma , a dedicated GM model calculator. Alternatively, we use the SPheno package Porod:2003um ; Porod:2011nf in combination with SARAH Staub:2013tta ; Staub:2012pb ; Staub:2010jh ; Staub:2009bi ; Staub:2008uz to compute the GM model spectrum.
We perform a comprehensive scan of the input parameters in Eq. (20) over the following ranges:
$\begin{split}-\sqrt{4\pi}&<{\lambda_{2,3,4,5}}<\sqrt{4\pi},\\ 120~\mathrm{GeV}<m_{h}^{\rm In}&<130~\mathrm{GeV},\quad 90~\mathrm{GeV}<m_{H}^{\rm In}<102~\mathrm{GeV},\\ 0<\sin{\theta}&<0.4,\quad\pi/4<|\alpha|<\pi/2.\end{split}$ (29)
To assess the global viability of the GM model in interpreting the observed excesses, we construct a likelihood-based framework. The total likelihood $\mathcal{L}_{\rm tot}$ is defined as
$\mathcal{L}_{\rm tot}\,=\,\exp\!\left(-\frac{1}{2}\,\chi^{2}_{\rm tot}\right)\times\Theta_{\rm PD}\,\Theta_{\rm unit}\,\Theta_{\tt HB}\,\Theta_{\tt HS},\qquad\chi^{2}_{\rm tot}=\chi^{2}_{\rm mass}+\chi^{2}_{95}+\chi^{2}_{\rm EW}+\chi^{2}_{\kappa}+\chi^{2}_{B}.$ (30)
The total $\chi^{2}_{\rm tot}$ contains only genuine likelihood contributions from the experimental observables, while the factors $\Theta_{\rm PD}$, $\Theta_{\rm unit}$, $\Theta_{\tt HB}$, and $\Theta_{\tt HS}$ act as acceptance criteria that restrict the likelihood support to theoretically and phenomenologically viable regions of parameter space.
The individual $\chi^{2}$ and $\Theta$ terms entering Eq. (30) are defined as follows.
  • $\chi^{2}_{\rm mass}=\left(\dfrac{m_{h}-125.18~\rm{GeV}}{1~\rm{GeV}}\right)^{2}+\left(\dfrac{m_{H}-95.4~\rm{GeV}}{1~\rm{GeV}}\right)^{2}$.
    Here, $m_{h}$ and $m_{H}$ are the theoretical predictions of the two $CP$-even Higgs states, and the total uncertainties (theoretical and experimental) are both set to $1~{\rm GeV}$.
  • $\Theta_{\tt HB}$ implements the direct-search constraints from HiggsBounds-5.10.0 Bahl:2022igd ; Bechtle:2008jh ; Bechtle:2011sb ; Bechtle:2013wla ; Bechtle:2020pkv . We define $\Theta_{\tt HB}=1$ for parameter points that satisfy the HiggsBounds exclusion test and $\Theta_{\tt HB}=0$ otherwise. In this way, direct-search constraints act as a veto on the parameter space rather than as a compensating contribution to the global $\chi^{2}$.
  • $\Theta_{\tt HS}$ encodes the consistency requirement from HiggsSignals-2.6.2 Bechtle:2013xfa ; Bechtle:2014ewa ; Bechtle:2020uwn ; Bahl:2022igd . To avoid double counting Higgs signal-strength information with the explicit $\chi^{2}_{\kappa}$ term defined below, we do not include the HiggsSignals output as an additive likelihood contribution in $\chi^{2}_{\rm tot}$. Instead, we require that the sample satisfy $\chi^{2}({\rm GM})<\chi^{2}({\rm SM})$, where $\chi^{2}({\rm SM})\approx 152.5$ denotes the SM reference value returned by HiggsSignals for the 125 GeV Higgs rate measurements in our setup. We therefore set $\Theta_{\tt HS}=1$ when this condition is fulfilled and $\Theta_{\tt HS}=0$ otherwise.
  • $\chi^{2}_{95}=\left(\dfrac{\mu_{\gamma\gamma}-0.24}{0.09}\right)^{2}+\left(\dfrac{\mu_{\tau\tau}-1.2}{0.5}\right)^{2}$: denotes the two $95~{\rm GeV}$ excesses, where $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ are the theoretical predictions including loop corrections.
  • The contribution $\chi^{2}_{\rm EW}$ encodes constraints from electroweak precision measurements in terms of the oblique parameters $S$, $T$, and $U$, which capture the one-loop contributions of the additional scalar states to gauge-boson self-energies and neutral-current processes Peskin:1990zt ; Peskin:1991sw . In this work, the $\chi^{2}_{\rm EW}$ is implemented using a multivariate Gaussian form,
    $\chi^{2}_{\rm EW}=\Delta\vec{O}_{\rm EW}^{\,T}\,\rho_{\rm EW}^{-1}\,\Delta\vec{O}_{\rm EW},$ (31)
    where
    $\Delta\vec{O}_{\rm EW}^{T}=\left(\frac{S+0.04}{0.10},\;\frac{T-0.01}{0.12},\;\frac{U-0.05}{0.09}\right)$ (32)
    and $\rho_{\rm EW}$ represents the correlation matrix, defined as
    $\rho_{\rm EW}=\begin{pmatrix}1&0.93&-0.70\\[4.0pt] 0.93&1&-0.87\\[4.0pt] -0.70&-0.87&1\end{pmatrix}$. (33)
    In addition, the $\rho$ parameter, which characterizes possible new sources of $\rm SU(2)$} breaking beyond the SM Higgs doublet, is constrained to be
    $\rho=1.00031\pm 0.00019.$ (34)
    In the custodial-symmetric GM setup, $\rho_{\rm tree}=1$ holds by construction, and any residual deviation arises only at loop level.
  • The contribution $\chi^{2}_{\kappa}$ encodes the constraints on the Higgs coupling modifiers through a two-dimensional Gaussian likelihood in the $(\kappa_{V},\kappa_{F})$ plane, and is defined as
    $\chi^{2}_{\kappa}=\Delta\vec{\kappa}^{\,T}\,V^{-1}\,\Delta\vec{\kappa},$ (35)
    where
    $\Delta\vec{\kappa}=\begin{pmatrix}\kappa_{V}-\hat{\kappa}_{V}\\ \kappa_{F}-\hat{\kappa}_{F}\end{pmatrix}$, (36)
    and $V$ denotes the corresponding covariance matrix,
    $V=\begin{pmatrix}\sigma_{V}^{2}&0.39\,\sigma_{V}\sigma_{F}\\ 0.39\,\sigma_{V}\sigma_{F}&\sigma_{F}^{2}\end{pmatrix}$. (37)
    In our analysis, we adopt the ATLAS combined fit results shown in Fig. 4 of Ref. ATLAS:2022vkf , with $\hat{\kappa}_{V}=1.035\pm 0.031$ and $\hat{\kappa}_{F}=0.95\pm 0.05$, and a correlation coefficient of $0.39$.
  • $\chi_{B}^{2}=\left(\dfrac{\mathcal{B}(B_{s}\to\mu^{+}\mu^{-})-3.1\times{10^{-9}}}{1.3\times{10^{-9}}}\right)^{2}+\left(\dfrac{\mathcal{B}(B\to X_{s}\gamma)-3.43\times{10^{-4}}}{0.44\times{10^{-4}}}\right)^{2}$ accounts for constraints from rare $B$-meson decays, which are strongly suppressed in the SM and therefore provide sensitive probes of physics beyond the SM, in particular extended Higgs sectors CMS:2014xfa .
In addition to the experimental limits mentioned above, theoretical considerations further constrain the parameter space through the Heaviside step functions $\Theta_{\rm unit}$ and $\Theta_{\rm PD}$ in Eq. (30). These functions act as strict vetoes, restricting the likelihood support to regions of parameter space that are theoretically consistent, where $\Theta(x)$ equals one for $x\geq 0$ and zero otherwise.
The $\Theta_{\rm unit}$ enforces perturbative unitarity of scalar $2\to 2$ scattering, requiring that the zeroth partial-wave amplitudes obey $|{\rm Re}\,a_{0}|<\tfrac{1}{2}$ (equivalently $|a_{0}|<1$). In the GM model this translates into bounds on linear combinations of quartic couplings Hartling:2014zca ; Chowdhury:2024mfu ; Aoki:2007ah , leading to the conditions
$\begin{split}\sqrt{(6\lambda_{1}-7\lambda_{3}-11\lambda_{2})^{2}+36\lambda_{4}^{2}}+\left|6\lambda_{1}+7\lambda_{3}+11\lambda_{2}\right|&<4\pi,\\ \sqrt{(2\lambda_{1}+\lambda_{3}-2\lambda_{2})^{2}+\lambda_{5}^{2}}+\left|2\lambda_{1}-\lambda_{3}+2\lambda_{2}\right|&<4\pi,\\ \left|2\lambda_{3}+\lambda_{2}\right|<\pi,\qquad\left|\lambda_{4}+\lambda_{5}\right|&<2\pi.\end{split}$ (38)
Therefore, $\Theta_{\rm unit}=1$ if all conditions in Eq. (38) are satisfied simultaneously.
The positive-definiteness requirement is imposed to ensure the stability of the electroweak vacuum, such that the scalar potential remains bounded from below for large field values. The quantum corrections can significantly modify the shape of the scalar potential, potentially leading to violations of the tree-level bounded-from-below (BFB) conditions Hartling:2014zca . Such effects may induce deeper vacua at large field values or render the potential unbounded, implying that the electroweak vacuum preserving custodial symmetry may not be the true ground state. This situation is reminiscent of the metastability of the SM electroweak vacuum for a $125~{\rm GeV}$ Higgs boson.
To address this issue, we impose the positive definiteness constraints Du:2024mry to guarantee the stability of the custodial vacuum by requiring positivity of certain subdeterminants of the quartic coupling matrices in the one-loop RG-improved potential. For simplicity, the one-loop RG-improved scalar potential is parameterized by the dimensionless couplings $\lambda$, $\rho_{1,2,\dotsc,5}$, and $\sigma_{1,2,3,4}$ of the most general ${\rm SU(3)}_{C}\times{\rm SU(2)}_{L}\times{\rm U(1)}_{Y}$ invariant GM model. [^2] Dimensional parameters are eliminated via the minimization conditions for the VEVs, and thus do not affect the large-field analysis relevant to positive definiteness. In the custodial-symmetric limit these couplings map onto the standard GM quartics $\lambda_{1,2,\dots,5}$ as
$\begin{split}\lambda&=4\lambda_{1},\\ \rho_{1}&=4\lambda_{2}+6\lambda_{3},\quad\rho_{2}=-4\lambda_{3},\quad\rho_{3}=2(\lambda_{2}+\lambda_{3}),\quad\rho_{4}=4\lambda_{2},\quad\rho_{5}=4\lambda_{3},\\ \sigma_{1}&=4\lambda_{4}-\lambda_{5},\quad\sigma_{2}=2\lambda_{5},\quad\sigma_{3}=2\lambda_{4},\quad\sigma_{4}=\sqrt{2}\lambda_{5}.\end{split}$ (39)
With these definitions and when the $\rho_{5}$ and $\sigma_{4}$ are small, the positive definiteness constraints then reduce to requiring that all leading principal minors of $\mathcal{V}$ are strictly positive,
$\det\mathcal{V}_{n}>0\qquad(n=1,2,\dotsc,5),$ (40)
where $\mathcal{V}$ is defined as
$\mathcal{V}=\begin{pmatrix}\frac{1}{4}\lambda&\frac{1}{8}(2\sigma_{1}+\sigma_{2})&\frac{1}{4}\sigma_{1}&\frac{1}{2}\sigma_{3}~&\frac{1}{4}\sigma_{3}\\[4.0pt] \frac{1}{8}(2\sigma_{1}+\sigma_{2})&\frac{1}{2}(2\rho_{1}+\rho_{2})&\frac{3}{2}\rho_{2}&\rho_{4}&\frac{1}{2}\rho_{4}\\[4.0pt] \frac{1}{4}\sigma_{1}&\frac{3}{2}\rho_{2}&\rho_{1}+\rho_{2}&\rho_{4}&\frac{1}{2}\rho_{4}\\[4.0pt] \frac{1}{4}\sigma_{3}&\rho_{4}&\rho_{4}&2\rho_{3}&\rho_{3}\\[4.0pt] \frac{1}{2}\sigma_{3}&\frac{1}{2}\rho_{4}&\frac{1}{2}\rho_{4}&\rho_{3}&\frac{1}{2}\rho_{3}\end{pmatrix}$, (41)
and $\mathcal{V}_{n}$ denotes the $n\times n$ upper-left submatrix of $\mathcal{V}$. Therefore, the $\Theta_{\rm PD}$ in Eq. (30) is defined as
$\Theta_{\rm PD}=\prod_{n}\Theta(\det\mathcal{V}_{n}).$ (42)
In practice, the full matrix is used numerically, while the analytic form shown here corresponds to the simplified limit.
It should be noted that the coupling parameters (such as $\lambda$, $\rho_{1,2,\dotsc,5}$, and $\sigma_{1,2,3,4}$, etc.) must be replaced by their renormalization group evolved values in regions of large field values. These evolved parameters can then be used to determine the vacuum structure of the GM model. We utilize the iterative method described in Du:2024mry to assess whether the parameter space that adequately explains the observed excess signals satisfies the positive definiteness constraints. The gauge coupling values at the EW scale employed in this study are derived from Antusch:2013jca . To ensure that the vacuum maintains custodial symmetry, the boundary conditions for the scalar quartic couplings at the EW scale, which are crucial for the RGEs, are selected in accordance with Eq. (39). To determine the scale dependence of the quartic couplings within the one-loop RG-improved tree-level scalar potential, it is essential to evolve the input parameters to an arbitrary scale, especially in regions of large field values, using the appropriate beta functions. In our detailed numerical analysis, to establish the positive definiteness constraints, we evolve the relevant quartic couplings and mass dimension parameters from the EW scale ($M_{Z}=91.2~{\rm GeV}$) up to the Planck scale ($M_{\rm Planck}\approx 1.12\times{10}^{19}~{\rm GeV}$) or to the lowest Landau pole scale, where any of the quartic couplings diverge.
To ensure a numerically reliable assessment of the GM model’s ability to simultaneously explain the two observed excesses, we conducted an extensive exploration of the parameter space using a combination of random scans, Markov Chain Monte Carlo (MCMC) techniques, and MultiNest sampling. These complementary methods provide both broad global coverage and efficient identification of phenomenologically relevant regions with higher likelihood. In total, we sampled approximately $\mathcal{O}(10^{8})$ points, providing sufficient statistical coverage to ensure that our conclusions are robust and unaffected by scanning artifacts. Convergence was confirmed by verifying the stability of the profile likelihood (PL) with additional sampling.
5 Numerical results
In this section, we present detailed results from our sampling. For clarity, we have filtered out all samples that fail the HiggsBounds or HiggsSignals consistency requirements, the positive-definiteness requirement, or the perturbative unitarity constraints; thus, all samples shown in the figures satisfy $\Theta_{\tt HB}=1$, $\Theta_{\tt HS}=1$, $\Theta_{\rm PD}=1$, and $\Theta_{\rm unit}=1$.
5.1 Fit results and benchmarks
Figure 1: The profile likelihood (PL) map projected onto the plane of signal strengths for di-photon and di-tau channels, $(\mu_{\gamma\gamma},\mu_{\tau\tau})$.
Figure 1 displays the PL distribution projected onto the $(\mu_{\gamma\gamma},\,\mu_{\tau\tau})$ plane for the $95~\mathrm{GeV}$ scalar in the GM model, after applying all theoretical and experimental constraints described in Sec. 4. The color scale indicates the PL normalized to its global maximum. A well-defined likelihood concentration is observed in our fit, indicating that the GM model is capable of simultaneously enhancing the di-photon and di-tau signal strengths relative to the SM expectations. The highest-likelihood region is centered around $\mu_{\gamma\gamma}\sim 0.25$ and $\mu_{\tau\tau}\sim 0.5$, demonstrating that the reported di-photon excess can be accommodated within the preferred parameter region, while the di-tau channel is enhanced to a moderate but non-negligible level.
Refer to caption Refer to caption Refer to caption
Figure 2: Left: The PL map on the $(\kappa_{F},\kappa_{V})$ plane. Middle: The PL $\mathcal{L}_{\kappa}$ on the $(\mu_{\gamma\gamma},\mu_{\tau\tau})$ plane. Right: The PL $\mathcal{L}_{\rm tot}$ without $\mathcal{L}_{\kappa}$ contribution on the $(\mu_{\gamma\gamma},\mu_{\tau\tau})$ plane.
In summary, the main difficulty in accommodating the 95 GeV $\gamma\gamma$ and $\tau\tau$ excesses within the GM model originates from the tension with the precision $\kappa$ fits of the 125 GeV Higgs couplings, which restrict the viable region to a thin and highly correlated slice of parameter space. As shown in Eq. (24), the signal strength $\mu_{\tau\tau}$ is directly related to $\kappa_{F}$. To illustrate this tension between the excesses and the Higgs-coupling data, Fig. 2 shows the PL distribution on the $(\kappa_{V},\kappa_{F})$ plane, the profile behavior of $\mathcal{L}_{\kappa}\equiv\exp\left(-\frac{1}{2}\chi^{2}_{\kappa}\right)$, and $\mathcal{L}_{\rm tot}/\mathcal{L}_{\kappa}$. The left panel of Fig. 2 shows a striking feature of this distribution: the emergence of a narrow, elongated band with a clear positive correlation between $\kappa_{F}$ and $\kappa_{V}$. This behavior can be directly understood from the analytic expressions in Eq. (22), which show that both coupling modifiers are controlled solely by the two mixing angles $\alpha$ and $\sin{\theta_{H}}$. $\kappa_{F}$ and $\kappa_{V}$ are strongly correlated through the underlying doublet-triplet mixing structure, and their allowed values populate only a restricted trajectory in the $(\kappa_{F},\kappa_{V})$ plane. As a consequence, the GM model can only access a limited "half-ellipse" structure observed in the PL map. In the parameter region preferred by the fit, an increase in $|\cos\alpha|$ simultaneously enhances both $\kappa_{F}$ and $\kappa_{V}$, leading to the positively sloped band observed in the figure. Moreover, a large $\kappa_{F}$ implies a sizable doublet component in the lighter 95 GeV state $H$, enhancing its gluon-fusion production cross section and leading to tension with direct searches encoded in HiggsBounds. These combined effects carve out the characteristic blank wedge in the distribution. As a consequence, in the middle panel of Fig. 2, the best $\mathcal{L}_{\kappa}$ is located in the region of $\mu_{\tau\tau}$ around 0.4. In the right panel of Fig. 2, the remaining contribution to $\mathcal{L}_{\rm tot}$ is larger in the region with higher $\mu_{\tau\tau}$.
Table 1: Benchmark points in the GM model. The $\chi^{2}_{\rm tot}$ is defined in Eq. (30). A checkmark (✓) and a cross (✗) indicate whether a given benchmark point passes or fails the corresponding constraint. All dimensional quantities are given in units of GeV.
BP1 BP2 BP3 BP1 BP2 BP3 BP1 BP2 BP3
$\lambda_{2}$ 0.8211-0.8211 0.6069-0.6069 $0.0698$ $m_{H}$ $96.8491$ $96.68$ $95.22$ $\mu_{\gamma\gamma}$ $0.2899$ $0.2368$ $0.2445$
$\lambda_{3}$ $0.8150$ $0.8336$ 0.8805-0.8805 $m_{h}$ $124.0331$ $125.2977$ $125.0191$ $\mu_{\tau\tau}$ $0.7423$ $0.5245$ $0.5445$
$\lambda_{4}$ $0.5232$ 0.0417-0.0417 $0.4704$ $m_{H_{3}}$ $194.5481$ $214.3225$ $222.6281$ $\chi^{2}_{\rm tot}$ $19.795$ $21.220$ $7.605$
$\lambda_{5}$ $0.2216$ $0.8814$ 0.0418-0.0418 $m_{H_{5}}$ $290.7393$ $340.3623$ $311.1941$ $\mathcal{B}(H_{5}^{\pm\pm}\!\to\!W^{\pm}W^{\pm})$ $0.1757$ $0.0032$ $0.1059$
$\alpha$ 0.8075-0.8075 $1.0548$ 1.0399-1.0399 $\kappa_{F}^{h}$ $0.8538$ $0.9328$ $0.9330$ $\mathcal{B}(H_{5}^{\pm\pm}\!\to\!H_{3}^{\pm}W^{\pm})$ $0.8243$ $0.9968$ $0.8941$
$\sin\theta_{H}$ $0.1309$ $0.0287$ $0.1246$ $\kappa_{V}^{h}$ $0.9296$ $0.9200$ $0.9836$ Unitarity
$M_{H}^{\rm in}$ $96.84$ $96.6780$ $95.2205$ $S$ 0.0832-0.0832 0.0930-0.0930 0.0850-0.0850 Positive definiteness
$M_{h}^{\rm in}$ $124.0331$ $125.2977$ $125.0191$ $T$ 0.0136-0.0136 0.0012-0.0012 0.0181-0.0181 HiggsSignals
$U$ $0.0007$ $0.0003$ $0.0004$ HiggsBounds
To highlight the characteristic features of this surviving region, we present three benchmark points in Table. 1. BP1 features a relatively large triplet admixture, with $\sin\theta_{H}=0.131$ and $\alpha=-0.808$. This results in a significant suppression of the SM-like Higgs coupling $\kappa_{F}$. Consequently, the 95 GeV scalar attains a sizable $\mu_{\tau\tau}=0.742$ and the highest diphoton rate among the three points, $\mu_{\gamma\gamma}=0.290$, leading to the smallest $\chi^{2}_{\rm 95}$. However, it does not satisfy the HiggsSignals constraints. BP2 represents the opposite limit with a very small triplet VEV, $\sin\theta_{H}=0.029$. The 125 GeV Higgs is therefore more SM-like. The reduced mixing suppresses the $\tau\tau$ signal strength to $\mu_{\tau\tau}=0.525$, while the diphoton rate, $\mu_{\gamma\gamma}=0.237$, is mainly maintained by charged-scalar loop effects. This benchmark satisfies both HiggsSignals and HiggsBounds. Compared to BP1, BP3 provides a better solution with $\sin\theta_{H}=0.125$ and $\alpha=-1.040$, for which $\kappa_{V}^{h}=0.984$ is remarkably close to experimental data. This point achieves a balanced configuration, $\mu_{\tau\tau}=0.545$ and $\mu_{\gamma\gamma}=0.245$, while remaining consistent with all theoretical and experimental constraints. It exemplifies the thin, correlated region identified in Fig. 2, where the 95 GeV excesses can be accommodated without significantly disturbing the 125 GeV Higgs couplings.
Figure 3: Similar to Fig. 1, but projected onto the GM model parameter planes.
5.2 Parameter space and phenomenology
Figure 3 displays the PL projected onto the eight GM model parameter planes. A clear pattern emerges: while the quartic couplings $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ span relatively broad ranges, the physically relevant mixing parameters are tightly constrained. In the two $\lambda_{i}$ panels, no pronounced minima appear, indicating that the quartic interactions are not individually strongly constrained. Instead, these couplings primarily support the scalar mass spectrum and enable viable loop contributions, without requiring fine-tuned correlations among themselves. In contrast, the $(M_{H}^{\rm in},\alpha)$ projection reveals a highly concentrated high-likelihood region centered at $(95~{\rm GeV},-\frac{\pi}{3})$, directly corresponding to the mass of the observed excess. The narrow range of the mixing angle $\alpha$ is consistent with the data, and large portions of the $\alpha$ parameter space are excluded, indicating that doublet–triplet mixing is tightly constrained. For values of $\alpha$ with identical magnitude as shown in the benchmark points BP2 and BP3, $\kappa_{F}$ remains unchanged, as expected. However, the coupling $\kappa_{V}$ associated with positive $\alpha$ is significantly smaller than that for negative $\alpha$, resulting in a pronounced deviation from experimental constraints. Moreover, $\kappa_{V}^{H}$ corresponding to positive $\alpha$ is substantially larger than its counterpart for negative $\alpha$, rendering this scenario unable to survive the additional Higgs-search constraints in the HiggsBounds package. This feature underlies the $\kappa$ tension observed in Fig. 2: while $\alpha$ at $\frac{\pi}{4}$ would better explain the excess, the overall fit only allows $\alpha$ near $\frac{\pi}{3}$. A similar pattern appears in the $(M_{h}^{\rm In},\sin\theta)$ plane. Overall, Fig. 3 demonstrates that the triplet admixture must remain moderate.
Figure 4: Similar to Fig. 1, but shown in the oblique-parameter $(T,S)$ plane and in the branching-ratio plane for the rare $B$-meson decays $B^{0}_{s}\to\mu^{+}\mu^{-}$ and $B\to X_{s}\gamma$.
The results for the branching ratios of rare $B$-meson decays and the EW precision observables $T$ and $S$ are shown in Fig. 4. These rare $B$-decay processes are highly suppressed in the SM and are dominated mainly by top-quark and $W$-boson loops. In the GM model, new charged scalars such as $H_{3}^{\pm}$, $H_{5}^{\pm}$, and $H_{5}^{\pm\pm}$ also participate in these loop diagrams. In our results, the two branching ratios are centered around the experimental central values and exhibit a strong negative correlation. This anti-correlation can be understood at the level of the effective weak Hamiltonian. Both observables are loop-induced flavor-changing neutral-current processes and are therefore sensitive to the charged-scalar sector of the GM model, in particular to the custodial triplet state $H_{3}^{\pm}$. The radiative decay $B\to X_{s}\gamma$ is governed by the electromagnetic dipole operator
$\mathcal{H}_{\rm eff}\supset-\frac{4G_{F}}{\sqrt{2}}V_{tb}V_{ts}^{*}C_{7}^{\rm eff}\mathcal{O}_{7},\quad\mathcal{O}_{7}=\frac{e}{16\pi^{2}}m_{b}(\bar{s}\sigma^{\mu\nu}P_{R}b)F_{\mu\nu}.$ (43)
The branching ratio scales approximately as $\mathcal{B}(B\to X_{s}\gamma)\propto|C_{7}^{\rm eff}|^{2}$ Buras:1998raa ; Misiak:2015xwa . In contrast, the rare decay $B_{s}^{0}\to\mu^{+}\mu^{-}$ is described by
$\mathcal{H}_{\rm eff}\supset-\frac{4G_{F}}{\sqrt{2}}V_{tb}V_{ts}^{*}\left(C_{10}\mathcal{O}_{10}+C_{S}\mathcal{O}_{S}+C_{P}\mathcal{O}_{P}\right),\quad\mathcal{O}_{10}=(\bar{s}\gamma^{\mu}P_{L}b)(\bar{\mu}\gamma_{\mu}\gamma_{5}\mu),$ (44)
where $\mathcal{O}_{S,P}$ denote scalar and pseudoscalar operators. The corresponding branching ratio depends predominantly on $|C_{10}|^{2}$ in the absence of large scalar contributions Bobeth:2013uxa ; Hermann:2013kca . In the GM model, the charged scalars $H_{3}^{\pm}$, $H_{5}^{\pm}$, and $H_{5}^{\pm\pm}$ modify both $C_{7}$ and $C_{10}$ at the one-loop level. However, the loop functions entering $\Delta C_{7}$ and $\Delta C_{10}$ differ in structure and interfere differently with the SM contributions. For the parameter region relevant to the 95 GeV interpretation, the charged-scalar contribution tends to shift $C_{7}$ and $C_{10}$ in opposite effective directions relative to their SM values, leading to the nearly linear anti-correlation observed in Fig. 4. Turning to the EW oblique measurements, the $T$ parameter is often considered a sensitive probe of custodial symmetry breaking. In the GM model, custodial symmetry is preserved at tree level, so the main contributions to $T$ arise from mass splittings within the triplet and quintuplet states. For all samples, custodial symmetry remains nearly exact, resulting in $T\approx 0$. However, for larger $v_{\Delta}$ or varying $\alpha$, mixing between the light scalar $H$ and the 125 GeV Higgs can shift the masses of certain scalars, causing mass differences within custodial multiplets and leading to deviations in $T$. The $S$ parameter reflects corrections to the $Z$ boson self-energy and its mixing, originating from new physics states. In the GM model, the $S$ parameter is sensitive to light charged scalars. When the light scalar $H$ has a larger triplet component, such as in the region $\alpha\sim-\pi/3$, loop effects are enhanced, resulting in more significant deviations in $S$. Conversely, when the 125 GeV Higgs boson is near the alignment limit, such as for $\alpha\sim\pi/3$ and a small triplet vacuum expectation value $v_{\Delta}$, the scalar sector is dominated by the doublet component, and mixing with the triplet states is suppressed. In this regime, the additional custodial multiplet scalars either become heavier or couple more weakly to the electroweak gauge bosons. As a result, their loop-induced contributions to the gauge-boson self-energies are reduced, leading to smaller corrections to the oblique parameter $S$.
Figure 5: Similar to Fig. 1, but showing the $(m_{H_{5}},m_{H_{3}})$ panel and the $m_{H_{5}^{\pm\pm}}$ versus $\sin\theta\times\sqrt{\mathcal{B}\left(H_{5}^{\pm\pm}\to W^{\pm}W^{\pm}\right)}$ panel. The red line indicates the ATLAS search ATLAS:2023sua limit on doubly charged $H_{5}^{\pm\pm}$ in $W^{\pm}W^{\pm}$ final states.
Finally, Fig. 5 displays the profiled likelihood projected onto the $(M_{H_{5}},\,M_{H_{3}})$ plane. A compressed spectrum is observed between the custodial fiveplet and triplet masses, with $\Delta m(m_{H_{5}},m_{H_{3}})<\mathcal{O}(100~{\rm GeV})$. In particular, the robust mass hierarchy $M_{H_{5}}>M_{H_{3}}$ implies that the doubly charged scalar $H_{5}^{\pm\pm}$ tends to decay predominantly into $H_{3}^{\pm}W^{\pm}$, rather than through the conventional $W^{\pm}W^{\pm}$ mode. As shown in Table 1, the branching ratio $\mathcal{B}(H_{5}^{\pm\pm}\to W^{\pm}W^{\pm})$ is typically suppressed to below the ten-percent level, and for the benchmark point BP2 it can even reach the per-mille level. This feature has important implications for collider phenomenology.
The right panel of Fig. 5 shows the plane of $m_{H_{5}^{\pm\pm}}$ versus $\sin\theta\,\sqrt{\mathcal{B}(H_{5}^{\pm\pm}\to W^{\pm}W^{\pm})}$, together with the published ATLAS limit ATLAS:2023sua derived for the simplified topology with dominant $H_{5}^{\pm\pm}\to W^{\pm}W^{\pm}$ decays. We emphasize that this comparison is only indicative. In the cascade-dominated regime relevant to our preferred parameter space, the actual signal topology is instead controlled by $H_{5}^{\pm\pm}\to H_{3}^{\pm}W^{\pm}$, and the resulting event kinematics and analysis efficiencies need not coincide with those assumed in the published same-sign diboson search. Therefore, the experimental contour in Fig. 5 should not be interpreted as a strict exclusion bound on the full GM scenario in this region.
Nevertheless, the figure still illustrates an important trend: once the branching ratio into $W^{\pm}W^{\pm}$ is strongly suppressed, the direct sensitivity of the existing same-sign diboson searches is correspondingly weakened. A dedicated recast of the relevant ATLAS and CMS analyses, explicitly incorporating the cascade topology $H_{5}^{\pm\pm}\to H_{3}^{\pm}W^{\pm}$, would be required to assess the true collider reach for this part of the parameter space.
Additionally, for the Drell–Yan process, which produces $H_{5}$ pairs via $\gamma$ and $Z$ exchange ATLAS:2021jol , the production cross section depends only on the gauge coupling and the Higgs mass $m_{H_{5}}$, and is independent of $v_{\Delta}$ as discussed in the experimental searches. In our work, $m_{H_{5}}$ ranges from several hundred GeV, where the cross section is relatively small, making it difficult for current experimental searches to exclude the surviving parameter space. The process $H_{5}^{\pm\pm}\to\ell^{\pm}\ell^{\pm}$ proceeds via a lepton number violation coupling between the triplet $\chi$ and lepton fields, conducted at LEP OPAL:2001luy . However, as pointed out in Chiang:2012cn , this decay channel becomes comparable to the weak gauge boson decay modes only when $v_{\Delta}\lesssim 10^{-4}~\text{GeV}$, making it negligible in the parameter region of interest for our study.
5.3 Additional checks
As discussed in Sec. 4, our main fit uses the $(\kappa_{V},\kappa_{F})$ likelihood rather than the full HiggsSignals output, in order to avoid double counting Higgs signal-strength information. It is nevertheless useful to verify that the preferred region identified in this reduced Higgs-likelihood treatment remains broadly consistent with the more detailed channel-based evaluation provided by HiggsSignals. The left panel of Fig. 6 shows the relation between $\chi^{2}_{\kappa}$ and $\chi^{2}_{\rm HS}$, with the color scale indicating the total profile-likelihood ratio.
Figure 6: Similar to Fig. 1. Left: relation between $\chi^{2}_{\kappa}$ and $\chi^{2}_{\rm HS}$, where $\chi^{2}_{\kappa}$ is derived from the $(\kappa_{V},\kappa_{F})$ plane analysis and $\chi^{2}_{\rm HS}$ is computed by the HiggsSignals package. Right: profile-likelihood distribution projected onto the signal strengths in the $\gamma\gamma$ and $b\bar{b}$ channels.
One can clearly see that the parameter points with small $\chi^{2}_{\kappa}$ also tend to yield comparatively small $\chi^{2}_{\rm HS}$ values. This result demonstrates that the simplified $(\kappa_{V},\kappa_{F})$ fit effectively captures the primary Higgs-coupling constraints relevant to our analysis, while HiggsSignals provides a more comprehensive channel-level evaluation. The spread visible in Fig. 6 indicates that the two approaches are not identical, as anticipated. In particular, the $(\kappa_{V},\kappa_{F})$ fit is based on a compressed description of the 125 GeV Higgs data and therefore does not retain the full channel-by-channel information, production-mode decomposition, or experimental correlations that are incorporated in HiggsSignals. For this reason, the $\kappa$-plane treatment should be regarded as an effective approximation rather than a complete replacement of the full Higgs likelihood. In practice, it is also somewhat more conservative than the full channel-based HiggsSignals evaluation, since the compressed $(\kappa_{V},\kappa_{F})$ description retains less experimental information and therefore has weaker discriminating power in excluding parameter points. Nevertheless, this comparison tests the robustness of our results. The conclusion that the $95~{\rm GeV}$ $\tau\tau$ signal intensity is constrained by $125~{\rm GeV}$ Higgs boson data holds true regardless of whether we use the simplified $\kappa$-plane method or the more detailed channel-based evaluation method.
The right panel of Fig. 6 shows a related feature of the surviving parameter space in the $b\bar{b}$ channel. Although the LEP $b\bar{b}$ excess is not included as an input constraint in our fit, some viable parameter points remain compatible with the corresponding preferred signal-strength range. In the GM model, the $b\bar{b}$ and $\tau\tau$ signal strengths are not independent, since both are controlled by the same underlying singlet–doublet mixing structure and share the common factor $|\kappa_{F}^{H}|^{2}\cdot\Gamma_{h}^{\rm SM}/\Gamma_{H}^{\rm tot}$, as discussed in Sec. 3. As a result, once the preferred region is selected by the combined $\gamma\gamma$ and $\tau\tau$ fit, the corresponding prediction for $\mu_{b\bar{b}}$ is automatically restricted to a relatively narrow range.
In particular, the best-fit region with $\mu_{\tau\tau}\sim 0.45$ tends to predict a $b\bar{b}$ signal strength close to the central value of the LEP excess. We do not interpret this as a primary fitting target of the present work, but rather as a correlated outcome of the surviving region selected by the LHC-driven channels. This behavior can be understood qualitatively from the fact that, unlike the $\tau\tau$ excess, the $b\bar{b}$ signal strength does not encounter the same direct tension with the precision data of the 125 GeV Higgs boson. Consequently, once the Higgs-coupling constraints compress the viable parameter space, the corresponding prediction in the $b\bar{b}$ channel can still remain close to the LEP preferred range.
Our result should not be viewed as being in conflict with earlier studies, but rather as a complementary perspective. Ref. Chen:2023bqr employed a more complete treatment of the Higgs-sector constraints and concluded that the GM model can account for the diphoton excess and remain compatible with the LEP $b\bar{b}$ excess, but cannot reproduce the large central value of the CMS $\tau\tau$ excess. Ref. Ahriche:2023wkj further showed that a larger $\tau\tau$ rate can arise either in a tiny single-peak region or, more efficiently, in a twin-peak setup involving multiple nearly degenerate states. In contrast, our analysis focuses on a single light $CP$-even state and shows that, within this setup, the $\tau\tau$ signal strength can be pushed to a moderate level of order $0.5$, while larger values are limited by the 125 GeV Higgs data. In this sense, the $b\bar{b}$ result shown in the right panel of Fig. 6 should be understood as a correlated implication of the surviving parameter space, rather than as evidence for a fully unified explanation of all three excesses.
6 Conclusions
In this work, we demonstrated that the GM model offers a unified explanation for the LHC hints near 95 GeV in the $\gamma\gamma$ and $\tau\tau$ channels using a single light $CP$-even custodial singlet. Implementing the RG-improved positive-definiteness conditions for vacuum stability enlarges the viable parameter space relative to traditional bounded-from-below requirements, while remaining consistent with perturbative unitarity, electroweak precision data, $B$-physics observables, direct searches, and Higgs signal-strength measurements.
Our fitting results clearly demonstrate the predictive characteristics of the parameter space: the singlet mixing parameter $\alpha$ is confined to a narrow interval around $\frac{\pi}{3}$; the triplet VEV $v_{\Delta}$ is at the few 10 GeV scale; and the custodial multiplets exhibit a mass hierarchy with $m_{H_{3}}<m_{H_{5}}$ at the ${\cal O}(10^{2})~{\rm GeV}$ scale. The $\gamma\gamma$ production rate is enhanced by charged-scalar loops, most notably from the doubly charged state $H_{5}^{\pm\pm}$, which interfere constructively with the SM contributions. The GM model can naturally accommodate the diphoton excess, but due to the tension from the Higgs data, the di-tau signal strength $\mu_{\tau\tau}$ can only be pushed to around 0.5, which is about $1\sigma$ away from the experimental central value.
These features make our framework highly testable in future experiments. At the HL-LHC, updated searches in the $\gamma\gamma$ and $\tau\tau$ final states, along with direct probes of $H_{5}^{\pm\pm}$, $H_{5}^{\pm}$, and $H_{3}^{\pm}$, will be able to explore much of the remaining parameter space. Future lepton colliders, such as CEPC and FCC-ee, will further offer decisive tests of this scenario.
Acknowledgements.
We thank Cheng-Wei Chiang and Martin J. White for helpful discussions. We acknowledge the referee’s insightful comments. This work was supported by the National Natural Science Foundation of China (Grant Nos. 12447167 and 12275067), by the Joint Fund of Henan Province Science and Technology R&\&D Program (Grant Nos.225200810030, 225200810092 and 225200810014), by the Startup Research Fund of Henan Academy of Sciences (Grant No. 231820011), by the Basic Research Fund of Henan Academy of Sciences (Grant No. 240620006), by the Graduate Innovation Fund of Henan Academy of Sciences (Grant No. 243320031), by the National Key R&\&D Program of China (Grant No. 2023YFA1606000) and by the Science and Technology Innovation Leading Talent Support Program of Henan Province (Grant No. 254000510039). PZ is supported by the ARC Discovery Project DP220100007 and by the Centre for the Subatomic Structure of Matter (CSSM).
Data Availability Statement
All figures presented in this work were generated using the Jarvis-PLOT framework. The complete plotting configuration files used in this study, together with additional result figures not shown in the main text, are publicly available on the webpage:
https://pengxuan-zhu-phys.github.io/Jarvis-Docs/jarvisplot/pub/gm95excess/